A Complete Bibliography of Publications in Numerische Mathematik (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

22 April 2024 Version 1.44

Title word cross-reference

1 [LW21]. 2 [BD21]. ^{k} [LLMP24]. B [DP20]. C^{0} [BySTZ21, LZ20]. E [Kri20]. ℓ^{1} [MH21]. h [LSW22, MCT20]. H(curl) [CW20]. H(div) [CW20]. H^{1} [CW20, FS20]. H^{2} [AP21]. J [AHP21]. L^{2} [HKP20]. L^{∞} [PR23]. Q [BNW20]. H [FMP22]. H^{2} [BBF22]. Ω [NP21b]. p [CFEV21, DFTW20, DHW23, Loi20, Zen23]. Q^{2} [LZ20]. QR [BDG20]. s [CIV21]. T [Hal21].

-bases [DP20]. -coercivity [Hal21]. -compatibility [Hal21]. -FEM [LSW22]. -Laplace [DHW23]. -Laplacian [Loi20]. -matrix [BBF22, FMP22]. -method [AHP21]. -Poisson [DFTW20]. -quasi-norm [Zen23]. -regularization [MH21]. -robust [CFEV21]. -schemes [Kri20]. -stable [NP21b]. -tensor [BNW20]. -term [CIV21]. -version [MCT20].

2D [DBR22, Ji23].

3D [Fun22, Ji23].

4 [Cia22].

accuracy [AHV24]. acoustic [CCWG^{+}24, CWHMB21, ER20, Lei22]. activation [LLMP24]. adapted [GJT20]. Adaptive [CP24, EMPS20, BC24, BBF22, BGIP23, CT22, CLX21, FPT20, HPSV21, LY22, MP21, PS22, Sei22]. adaptivity [MP23]. adjoint [BKR21]. ADMM [GSY20]. advection [CHHM22]. Alfeld [CGGH24]. algebraic [BGJR23, HPSV21]. algebraically [JK22]. algorithm [BC24, Ber20, CLL^{+}22, FP22, HLW20, KLL21, LO20, NSD20, LO21]. algorithms [BAMR22, HJK21, HPSV21, HHMX21, LL21b, Loi20]. Allen [CP23]. alternating [HK22b]. Ampère [Ber20, BySTZ21, GT24]. Analysis [CP23, HMW23, Liu21, XY21, BB22, Ber21, BCD20, BMS23b, BDK^{+}20, BDG22, CKP22, CNRS23, CJT21, CHL21a, Cia22, CZ21, CK20, EG23, EGK22, FRZ21, GQS23, GF22, HL20a, HK22b, HHMX21, JWL22, Lei22, MSSCT21, MP21, MP23, Nat20a, Nat20b, OS22, SS23a, SSVS21, SMS20, Sun22]. analytical [AA21]. anisotropic [BL24, Kop20]. ANOVA [LPU23]. anti [dAFR20]. anti-Gauss [dAFR20]. any [BGM22a]. application [BCG20, DJC20, HL21a, HL21b, KKK^{+}22a, Lei22, RLK20, XY21]. applications [JNPS23, LL21b, WZL22]. applied [Le 23]. approach [ALP23, CDE20, DES21, EF21, FLMM20, GSY20, JL20, JK20, MSSCT21]. approaches [CHHM22]. approximating [Tro22]. Approximation [MRT20, APR22, AB21, AK22, ABNP21, BCG20, BD21, BDK^{+}20, CMR21, CIV21, Cia22, DJC20, FHS20, GN21, Hal21, HK22a, KKK^{+}22a, KNV21, KW23, KN21, LL21a, LLMP24, MCT20, MNR22, MH21, NP21a, OS22, SMS20, SSW20, Sei22, TNW21, Wan21, ZJN21]. approximations [AHV24, AN20, BBF22, CCZ20, CP23, FMP22, GT24, HKP20, JNPS23, MSSCT21, SS22]. arbitrary [LL21a, SMS20]. artificial [HL22]. assimilation [BGM22b]. Asymptotic [BMS23b, KLMNS22, CHY23, EG23, FRZ21, HK22a, LSV21]. asymptotic-preserving [CHY23, LSV21]. asymptotics [XL20]. Audusse [GJT20]. augmented [BCD20]. augmented-wave [BCD20]. autonomous [BOS22]. auxiliary [CMS20]. averaged [DHW23]. aware [HI23]. axisymmetric [BBMRB21].

back [JL20]. back-and-forth [JL20]. backward [BPR21]. Bakushinkii [HJP20]. balanced [FLNS21]. Banach [Rea24]. barotropic [BCG20]. Barron [LLMP24]. barycentric [FCH22a, FCH22b]. based [BGJR23, BHL22, CW20, CZ21, HJK21, Ji23, KKK^{+}22a, KW23, MSSCT21, MP21, SSW20, Sei22]. bases [CIV21, DP20]. basis [AK22, DS22, Pag21]. BayesCG [RICO23]. Bayesian [HK22a, SSW20]. BDDC [Huy22]. behaviour [CHHM22]. Bellman [BPR21, CCDS23]. Beltrami [Guo20]. BEM [FMP22]. bending [BR20, FHS20]. bending-torsion [BR20].

Bernstein [AK22]. best [CIV21, Wan21]. Beyond [HJP20]. bidomain [Huy22]. bilinear [BGW23]. Bingham [CS23a, GL22]. Biot [CLY22]. block [AK20, Tro22]. blocks [AK20]. blow [MT20]. blow-up [MT20]. Boltzmann [BGM22a]. Boris [HLW20]. bound [BL22, LSW22]. Boundary [CWHMB21, Sei22, ADGS23, CCWG^{+}24, CCDS23, DLO20, EBL22, EOR22, GN21, GL20, GÖSS20, GSUT21, HK23, HL21a, HL21b, HSS^{+}20, JK20, KK20, KH20, KPP^{+}21, KN21, Lei22, MCT20, Mat22, MP23, Nab21, SMS22, WM21, XY21]. bounded [CIV21, Li21]. Bounds [AK22, CEP21, CP24, GLW21, HI23, LV22]. Bounds-constrained [AK22]. Boussinesq [XRV21]. Brillouin [CEG^{+}20]. Brinkman [BBMRB21]. Brockett [BBFR21]. Budget [XN23]. Budget-limited [XN23]. Bulk [FMS21, CDE20]. Bulk-surface [FMS21]. BV [GJT20].

Caccioppoli [FMP22]. Caccioppoli-type [FMP22]. Cahn [ABNP21, BK22b, CP23, Nab21]. calculations [HMW23]. Carlo [LRS20, GV23, GKK^{+}24, LRS19, LSV21, SSW20]. cells [Ber21]. change [KPW20]. changing [Cia22]. characteristic [CS23b]. charged [HLW20, HL20a, HLS22]. charged-particle [HLW20, HL20a, HLS22]. chemistry [BDP22]. Chorin [KS20]. class [HK22b, KLMNS22]. closed [EGK22, KLL21]. cluster [HMW23]. clustering [TNW21]. clusters [DBR22]. coarse [Nat20a, Nat20b]. coefficient [Cia22]. coefficients [EMPS20, JLZ20, NN21]. coercive [CMR21]. Coercivity [CWS22, CWS23, DHL20, Hal21]. Cohen [CHL21a]. collocation [DP20, GSH23]. compactifications [MT20]. compatibility [Hal21]. complement [SNO20]. complements [DBR22]. Complete [HK23]. completion [ZJN21]. complex [CGGH24, TLO22a, TLO22b]. complexity [BGIP23, MMRV24]. compressible [FLMS23, KLMNS22, KN21, MNR22, SS22, Sun22]. computable [LV22]. Computation [BRS23, ERSS20, KH20]. computational [BGIP23, BKR21]. computations [Uhl24]. Computing [MS20, BAMR22, GSH23]. Concentration [CHY23]. condition [Mat22, XY21, ZJK23]. conditions [CCDS23, DLO20, EOR22, HK23, HL21b, JK20, KH20, KN21, Lei22, Nab21, ON22, HL21a]. cones [CW20]. conforming [AP21, CW20, GK20, ZZ22]. consensus [HJK21]. consensus-based [HJK21]. conservation [AHV24, GJT20, GTV22, GL20]. Conservative [CJT21]. constrained [AK22, BNW20, GSTU20, GSY20, GKK^{+}24, Jin22]. constraint [CZ22, WS21]. constraints [GF22]. Construction [KN21, AB21, BBF22]. Continuous [BGM22b, MS20]. continuous-time [MS20]. continuum [GH21]. control [AN20, BBFR21, BGW23, BCGT23, CMR21, CP23, FK23, GT24, GSY20, GKK^{+}24, HSS^{+}20, PK22, VZ22, Win20]. controls [MRT20]. convection [BNO20, BNO22, CS23b, JK22, Liu21]. convection-diffusion [BNO20]. convection-diffusion-reaction [JK22]. convection-dominated [BNO22]. converge [Wan21]. Convergence [BD21, CFS20, CS20, FPT20, FRZ21, GJT20, GGS22, HJK21, HPSV21, Jin22, BMS23a, BPR21, CP24, CK20, DS21, FS20, GF22, GL20, GSTU20, HT22, Har21, HLO20, JWL22, KS20, PS22, Sag22, SS23a, SSW20]. Convergent [CT22, CZ22, KL17, KLL21, NKL21]. convex [ATW22, BySTZ21, CT22, OS22]. convexity [BySTZ21]. convolution [JK20, MR21]. Corners [DNT21, HK23]. Corrected [WM21]. Correction [CWS23, CH21, FCH22b, HL21a, HW21a, LO21, LRS20, Nat20a, NKL21, OZ23, TLO22a, LL21b]. COS [Jun24]. cost [HPSV21]. could [Fun22]. coupled [BB22, BDG22, HMW23, WZL22, ZJK23]. coupling [BKR21, DS21, KL17, NKL21, ON22]. couplings [FMP22]. CP [ZJN21]. crack [CZB22]. cross [CFS20, CP22, Sei22]. cross-diffusion [CFS20]. cross-points [CP22]. Crouzeix [BOS22, BW21, Guo20, Ji23]. crystal [CLL^{+}22]. crystals [BNW20]. cube [NP21a]. cubed [BBC23]. cubic [LW21]. curvature [EGK22, GN21, HL22]. curves [AA21]. cuspidal [AL20]. cut [BHL20]. CutFEM [BHL22].

D [BD21, LW21]. Darcy [BB22]. data [BGM22b, DES21, GH21]. decay [DHL20, PS21, XL20]. decomposition [BS21, BMS23b, DNT21, DJC20, DBR22, GGS22, Nat20a, Nat20b, NN21]. deep [BTU22]. definite [FLNS21]. degenerating [KKK22b]. degree [LL21a]. Dekker [LO21, LO20]. deluxe [Huy22]. densities [Sag22]. density [LLW24]. dependent [BFM22, CS23a, GSH23, JLZ20, NKL22]. derivation [CFEV21, CJT21]. derivative [LLMP24]. derivatives [CCDS23, Win20]. deterministic [HI23]. develop [Fun22]. Development [CHL21a]. DG [MSSCT21]. Difference [BGM22a, HH20, LZ20]. differential [BGJR23]. differential-algebraic [BGJR23]. differentiation [BPR21]. diffusion [BMS23a, BNO20, BNO22, CCZ20, CFS20, CHHM22, CS23b, DS22, EGK22, FLNS21, JK22, Liu21, NN21, Rie23]. Diffusion-dominated [BNO20]. diffusive [LSV21]. dilation [CWHPV23]. dimensional [AN20, CGGH24, CH20, CH21, CZB22, GH21, JZZ22, KW23, Liu21, MP23, Yse20]. dimensions [FMS21, WM21, Yse24]. Direct [ATW22, ABMS21, Liu21, SNO20]. direction [HK22b]. Dirichlet [HSS^{+}20, Win20, XY21]. Dirichlet-to-Neumann [XY21]. Discontinuous [AN20, ABMS21, BBJP20, CS23b, DRG^{+}20, GK20, GLW21, GTV22, HK22b, HHMX21, HSS^{+}20, Liu21, WS21]. discrete [BDG22, CGGH24, DHL20, ERSS20, GLW21, GSTU20, HJK21, HKP20, Har21, HK22b, HH20, Ji23, KL17, Kri20, Kri21, LZ20, LW21, NKL21, SMdD22, XRV21]. discretisations [CS23a]. discretization [DHW23, HL21a, HL21b, HK22b, PK22]. discretizations [Huy22]. discretized [Uhl24]. dispersive [CJT21, EOR22]. displacement [ABMS21, BGM22b, Sun22]. dissipation [CZ22]. dissipative [HH20, SX21]. distance [MS20]. distributed [FK23]. distribution [HJP20, XN23]. div [ZZ22]. div-conforming [ZZ22]. Divergence [GK20]. Divergence-conforming [GK20]. does [Wan21]. Domain [BS21, ADGS23, AL20, BMS23b, CDE20, DNT21, DBR22, GÖSS20, GGS22, KK20, KH20, Nat20a, Nat20b]. domains [AA24, BDP22, BySTZ21, BFM22, CWS22, CWS23, CWHPV23, CDE20, DS21, HK23, HHK^{+}24]. dominated [BNO20, BNO22]. Double [Rie23, CWHPV23]. double-layer [CWHPV23]. DPG [FHS20, FHN22]. dqds [FP22]. driven [CZ22]. DtN [LY22]. dual [BK22a, Jin22]. dynamic [BGJR23, HL21a, HL21b, Nab21]. Dynamical [Pag21, GQS23, KNV21]. dynamics [BTU22, HLW20, HL20a, HLS22, Pöt23].

E-schemes [Kri21]. edge [Cia22, DBR22]. effect [AJH21]. Efficient [Loi20, NP21a, BKR21, CZ21, KPW20]. eigenfunctions [LV22]. eigenspaces [GSH23]. eigenvalue [AHP21, AL20, CP24, FP22, GK20, Hal21, RDYL21]. eigenvalues [CEP21, Tro22]. elastic [ABMS21, BR20, DBR22, GN21, LY22]. elasticity [CGGH24, HHMX21]. elastodynamics [ADGS23, DRG^{+}20]. electrical [GH21]. electrode [GH21]. electromagnetic [Cia22, NKL22]. electronic [HMW23]. element [AJH21, BOS22, BL24, BW21, BGIP23, Ber21, BHL20, BNO20, BHL22, BNO22, CCWG^{+}24, CCZ20, CLY22, CWHMB21, CMS20, CLX21, CHL21a, CH20, CH21, CZB22, DST22, EBL22, ER20, FHKP24, FMS21, GÖSS20, Guo20, HKP20, HL22, JWL22, Ji23, KK20, KKK22b, KLL21, KPP^{+}21, LZ20, LY22, MCT20, NN21, PK22, Sei22, XY21]. elements [ADGS23, AP21, ATW22, CW20, Cia22, FK23, GK20, GL22, HW21a, HW21b, ZZ22]. elliptic [CEP21, CKP22, CNP22, CMR21, CLX21, CDE20, Har21, HL20b, HW21a, HW21b, HSS^{+}20, MP23, NN21, SSVS21, SNO20]. Energy [MNR22, BL24, CH20, CH21, EG23, HH20, LCO24, SX21, Win20]. energy-dissipative [HH20]. enforcing [BySTZ21]. enriched [HW21a, HW21b]. ensemble [BBFR21, CHL^{+}21b, PS22]. entropic [GKK^{+}24]. Entropy [GF22, Kri20, Kri21, BRS23, BS21, BCG20, GJT20]. enumeration [Chk24]. Equation [BKR21, ABNP21, BBJP20, BySTZ21, BNO20, BNO22, CFEV21, CP23, GT24, GGS22, Guo20, HK23, JZZ22, KH20, Le 23, LM21, LW21, MR21, MRT20, Nab21, PK22, SS23a, Sei22, XY21]. Equation-free [BKR21]. equations [dAFR20, AA21, AN20, BBMRB21, Ber20, BK22b, BPR21, BCG20, BD21, BFM22, CHY23, CCDS23, CMR21, CHHM22, CWS22, CWS23, CDE20, DHL20, FPT20, FLNS21, Fun22, HMW23, HL21a, HL21b, HH20, KNV21, KL17, KLMNS22, KN21, Lei22, Li21, Liu21, Mat22, Pöt23, SX21, TLO22a, TLO22b, WM21, WS21, Yse20, Yse24, NKL21]. equilibrated [CCZ20]. Error [BK22a, BTU22, HK22b, Win20, XRV21, AM22, BL22, BCGT23, BK22b, BDG22, CCZ20, CKP22, CFEV21, CMS20, CH20, CH21, FLMS23, GT24, GLW21, GTV22, GÖSS20, HJK21, HKP20, HI23, HK22a, HL20b, KK20, Kop20, KV21, KPP^{+}21, LSW22, Lei22, LV22, MP21, MP23, Nab21, SSVS21, SMS20, SMS22]. essential [CWS22, CWS23]. estimate [GÖSS20, Nab21]. estimates [AM22, BK22a, BK22b, BTU22, CFEV21, CMS20, CH20, CH21, EG23, FMP22, FLMS23, GTV22, HJK21, HKP20, HK22a, HL20b, KK20, Kop20, KPP^{+}21, SMS20, SMS22, Win20, XRV21]. estimation [KV21, SNO20]. estimator [CCZ20]. estimators [BRS23]. Euler [BCG20, FLMM20, KLMNS22, LM21]. Eulerian [BFM22].

European [Jun24]. evolution [CLL^{+}22]. evolutionary [FPT20]. Evolving [HL22, BK22b, KLL21]. Exact [SMdD22]. existence [Li21, SNO20]. expansion [Yse20]. expansions [XL20]. Explicit [SMS20, HL21a, HL21b, SMS22]. Exponential [BMS23a, FS20, TNW21, CS23b, DHL20, Rie23]. extended [BHL22]. exterior [AA24, KL17, NKL21]. Extremal [DP20]. extreme [Tro22].

families [ZZ22]. Fast [BDG20, KKK^{+}22a, LPU23, BC24, JL20]. faster [Wan21]. FastTwoSum [LO21, LO20]. FEM [BMS23a, BNW20, CNRS23, EMPS20, FMP22, FS20, LSW22, NN21]. FEM-BEM [FMP22]. FEMs [CNP22, GQS23, Sun22]. FFT [CZ21]. Fiber [RDYL21]. field [BB22, HLW20, HL20a]. field-of-values [BB22]. fields [Cia22, CK20]. filling [CZ22]. filtered [FPT20, HLW20]. filtering [BC24, CHL^{+}21b, CZ21]. filters [BBMRB21]. Finite [BGM22a, CH20, HKP20, AHV24, AP21, ATW22, AJH21, BOS22, BL24, BW21, BGIP23, Ber21, BHL20, BNO20, BHL22, BNO22, CCZ20, CGT20, CZ22, CFS20, CLY22, CHHM22, CMS20, CLX21, CHL21a, Cia22, CZB22, ER20, FLMM20, FLMS23, FHKP24, FK23, GF22, GK20, GL20, GL22, HW21a, HW21b, HH20, HL22, JWL22, Ji23, KK20, KKK22b, KLL21, LZ20, LY22, Nab21, NN21, PK22, XY21, ZZ22, CH21]. finite-element [CMS20]. finite-volume [Nab21]. First [Füh20, DRG^{+}20, FPT20]. First-order [Füh20]. Fisher [BBJP20]. fixed [SS23a]. floating [HI23]. flow [EGK22, GN21, GL22, HL22, KLL21, ZJK23]. flows [BL24, BCG20, CGT20, CMS20, GSTU20, MNR22]. fluid [BTU22, BDG22, SS22, ZJK23]. fluid-structure [BDG22, SS22]. fluids [OS22, SS22]. flux [CLY22, GJT20, GTV22]. Fokker [JZZ22, DHL20, NSD20]. forcing [JZZ22]. form [AK20, DRG^{+}20]. Formulation [GL20, BGM22a, Ber21, FHS20, SNO20]. forth [JL20]. Fortin [DST22]. forward [Sag22]. FOSLS [GS24]. four [CNRS23]. Fourier [XY21]. fourth [CNRS23]. fourth-order [CNRS23]. fractal [CCWG^{+}24, CWHMB21, JK20]. fractional [BMS23a, DS22, FLNS21, NSD20, Rie23]. fractured [BHL20]. fragmentation [SS23a]. framework [AA21, BGW23, HT22, VZ22]. Fredholm [dAFR20]. free [BCGT23, BKR21, JWL22, Le 23]. frequency [EBL22, LSW22]. friction [ZJK23]. friction-type [ZJK23]. full [HMW23]. full-coupled [HMW23]. Fully [BDG22, LV22, GLW21, HK22b, HH20, Huy22, KL17, Kri20, Kri21, LW21, NKL21, XRV21]. fully-discrete [GLW21, Kri20, Kri21]. function [LLMP24]. Functional [KPP^{+}21]. functionals [APR22, MMV22]. Functions [PS21, APR22, Ji23, XL20, Yse20]. fundamental [AA24].

Galerkin [AN20, BBJP20, CWS22, CWS23, CS23b, DRG^{+}20, EBL22, EMPS20, GK20, GLW21, Hal21, HK22b, HHMX21, HSS^{+}20, Liu21, PK22, WS21]. gap [BOS22]. gas [EG23]. Gauss [APR22, dAFR20]. Gaussian [BL22]. GenEO [MCT20]. General [RLK20, BJPV20, KN21]. generalized [CK20, CZB22, FHKP24, IGS21]. generating [Van23]. generic [AB21]. geometric [BL24, Ber20]. Gevrey [FS20]. global [Har21]. Goal [BGIP23]. Goal-oriented [BGIP23]. Godunov [GJT20, GTV22]. grad [ZZ22]. graded [ADGS23]. gradient [CGT20, CMS20, CLL^{+}22, Jin22]. Grid [Le 23]. Grid-free [Le 23]. grids [AB21, KW23]. Gross [AHP21]. group [CS20]. Guaranteed [CEP21, CP24, CFEV21, GT24]. guarantees [CIV21].

Hamilton [BPR21, CCDS23, FPT20]. Hamiltonian [BGJR23, Pag21]. Hanke [Rea24]. hard [EBL22]. harmonic [CIV21]. harmonics [BBC23]. Hausdorff [CCWG^{+}24]. Hausdorff-measure [CCWG^{+}24]. HDG [SSVS21]. heat [TLO22a, TLO22b]. Helmholtz [AA21, CFEV21, DNT21, GGS22, HK23, LSW22, XY21]. HHO [BCGT23]. hierarchical [EMPS20, Sei22]. High [KW23, AP21, AB21, CEP21, CLX21, EBL22, JZZ22, LSW22, MP23, Yse24]. High-dimensional [KW23, MP23]. high-frequency [EBL22]. high-order [CEP21, CLX21]. Higher [ADGS23, CCZ20, CT22, NN21, Yse20]. Higher-order [ADGS23, CT22, NN21]. Hilliard [ABNP21, BK22b, Nab21]. Hilliard-type [BK22b]. Hölder [CK20]. holomorphic [Hal21]. homogeneous [MT20]. homogenisation [BKR21]. homotopy [RDYL21]. Hood [DST22]. Hopf [KS20]. hp [GLW21, ADGS23, BMS23a, FS20]. hp-discontinuous [GLW21]. hp-FEM [FS20]. Hybrid [LRS19, BJPV20, CEP21, CT22, CHHM22, DBR22, PK22, LRS20]. hybridizable [HSS^{+}20]. hyperbolic [LPU23, ON22]. hypersingular [MCT20]. hypocoercivity [DHL20].

II [BNO22, CIV21, HK23]. IIA [AM22]. ill [Jin22, PS22]. ill-posed [Jin22, PS22]. illposed [PR23]. image [JNPS23]. immersed [JWL22, Ji23]. impedance [GH21]. implementation [LZ20, ON22]. implementations [CZ21]. implicit [HL21a, HL21b, HK22b, Huy22, KLMNS22, LM21]. implicit-explicit [HL21a, HL21b]. implies [Li21]. Improved [FLMS23, GS24, KMZ23, JLZ20]. improvements [AK21]. including [HPSV21, Mat22]. Incomplete [JZ20]. incompressible [BDG22, GQS23, MSSCT21]. indefinite [CKP22, CNP22]. independent [BBF22]. inequalities [Kri20, Kri21]. inexact [Nat20a, Nat20b]. inflow [KN21]. inflow/outflow [KN21]. information [ABMS21]. initial [RLK20]. initial-value [RLK20]. inspired [ES24, FLMM20]. instability [MS20]. instationary [EG23]. integral [dAFR20, AA21, CWS22, CWS23, WM21]. integration [GKK^{+}24, KPW20]. integrator [HL20a, LW21]. integrators [CS20, HLS22, MV20]. integrodifference [Pöt23]. interacting [CFS20]. interaction [BDG22, EGK22]. interactions [SS22]. interface [ABNP21, CLX21, HW21a, HW21b, ZJK23]. interior [BySTZ21, KL17, NKL21]. internal [ABMS21, NSD20]. Interpolation [BBC23, SS23b, BGW23, FCH22a, FCH22b, IGS21, KKK^{+}22a, KMZ23, SMS22]. introduction [Uhl24]. invariant [CWHPV23]. inverse [BNO20, BNO22, DES21, HJP20, Har21, HK22a, SSW20, SNO20]. inverses [FMP22]. inversion [ABMS21, PS22]. investigation [BBFR21]. involving [APR22, SS22]. ion [Ber21, GF22]. isogeometric [BDK^{+}20, DS21, SMS20]. isotropic [FS20]. iteration [BGJR23, DFTW20, DJC20, Rea24]. iterations [BDG20]. Iterative [MMRV24, BC24, CZ21, HPSV21, JZ20, PR23, Yse24].

Jacobi [BPR21, CCDS23, FPT20]. joined [DBR22].

Kacanov [DFTW20]. Kalman [CHL^{+}21b, PS22]. Kernel [BBF22, KKK^{+}22a, KW23]. kernel-based [KKK^{+}22a, KW23]. Kernel-independent [BBF22]. kind [dAFR20, CWS22, CWS23]. knowing [HJP20]. KPP [BBJP20]. Krylov [AK21, BRS23, Huy22, PS21, RICO23, Tro22]. Krylov-BDDC [Huy22]. Kutta [MR21].

Lack [ALP23]. Lagrangian [CCDS23, SMdD22]. Landweber [Rea24]. Laplace [AA21, DHW23, Guo20, HK22a, SSW20, Yse20, Yse24]. Laplace-based [SSW20]. Laplace-like [Yse20, Yse24]. Laplacian [Loi20]. Large [HLS22, BRS23, MP21]. Large-stepsize [HLS22]. Lattice [Chk24, BGM22a, KKK^{+}22a]. lattice-point [KKK^{+}22a]. Lavrentiev [BOS22]. laws [AHV24, GJT20, GTV22, GL20]. Lawson [HLO20]. layer [CWHPV23, DRG^{+}20, LLMP24]. layers [CS23b, NKL22]. leak [ZJK23]. learning [BTU22, XN23]. Least [FK23, Füh20, LM21]. Least-squares [FK23, Füh20, LM21]. Legendre [Wan21]. Leray [KS20]. level [Ji23, MCT20, Nat20a, Nat20b, SSW20]. Lie [CS20]. like [DJC20, Yse20, Yse24]. limit [ABNP21, LSV21]. limited [XL20, XN23]. linear [ABMS21, ALP23, BGJR23, BK22b, CKP22, CNP22, CHHM22, Chk24, EOR22, FCH22a, FCH22b, GSY20, HJP20, HK22b, HHMX21, Jin22, KK20, MMV22, PS22, PR23, SSVS21, VZ22]. linear-quadratic [VZ22]. linearization [HPSV21, Zen23]. linearized [SNO20]. linearly [KLMNS22]. Lipschitz [CWS22, CWS23, CWHPV23]. liquid [BNW20]. Lithium [Ber21]. Lithium-ion [Ber21]. local [BC24, CS23b, LLW24]. Locality [LLW24]. localized [FHKP24]. locally [CWHPV23]. locally-dilation-invariant [CWHPV23]. lognormal [EMPS20]. Long [CHHM22, HL20a]. Long-term [HL20a]. Long-time [CHHM22]. loosely [BDG22]. Lotka [CHY23]. low [BDG20, BCG20, BV22, JNPS23, KNV21, KMZ23, LW21, MMRV24]. low-rank [BV22, KMZ23]. low-regularity [LW21]. Lower [Kop20, CEP21, CP24]. lowest [CKP22]. lowest-order [CKP22]. lumped [ER20].

MAC [FLMS23]. Mach [BCG20]. magnetic [HLW20, HL20a]. magnetized [FRZ21]. magnetohydrodynamics [GQS23]. makes [MV20]. manifolds [GN21]. mapping [CDE20]. mass [ER20]. mass-lumped [ER20]. matched [DRG^{+}20]. matching [LL21b]. Matérn [CK20]. Mathematical [Nat20b, Nat20a]. matrices [AK21, BRS23, BDG20, DP20, PS21, Sei22]. matrix [APR22, BBF22, BV22, FMP22, MS20, MSSCT21, NP21b, Uhl24]. matrix-theoretic [MSSCT21]. Maximal [MH21, AM22]. Maximizing [WZL22]. maximum [BDP22, LZ20]. Maxwell [NKL21, AJH21, KL17]. MDFEM [NN21]. mean [EGK22, HL22]. means [dAFR20]. measure [CCWG^{+}24, ERSS20, GKK^{+}24]. measurements [GH21]. measures [LRS19, LRS20]. mechanics [SMdD22]. media [BHL20, Sun22]. medium [BGM22b]. meets [LPU23]. mesh [Li21]. meshes [ADGS23, BJPV20, Kop20]. method [AHP21, AA24, AK21, BOS22, BL24, BC24, BJPV20, BL22, BCD20, BySTZ21, BHL20, BNO20, BNO22, CCWG^{+}24, CEP21, CLY22, CWS22, CWS23, CLX21, CHL21a, CS23b, CZB22, DS22, DS21, DRG^{+}20, ER20, EOR22, FHKP24, FMS21, Füh20, FHN22, GL20, GÖSS20, GSH23, HMW23, HW21a, HW21b, HSS^{+}20, JL20, Ji23, Jin22, Jun24, KK20, KS20, LLS21, Le 23, LM21, LZ20, LY22, LSV21, Mat22, MP21, NN21, RDYL21, SSVS21, Yse24, ZJN21]. methods [AM22, BGIP23, BRS23, BTU22, CWHMB21, CP22, DNT21, EBL22, ES24, GLW21, GGS22, HLO20, HL22, KW23, KPP^{+}21, LRS19, LRS20, Liu21, Nat20a, Nat20b, OZ17, OZ23, Pag21, RLK20, SSW20, Sei22, Tro22, WS21, XY21, Zen23, Zha21, ZJK23]. metric [ERSS20]. MHD [WS21]. Mimicking [GH21]. Mindlin [FHS20]. minimization [BK22a, CT22, Zen23]. Minimum [LL21b, LCO24]. Minimum-correction [LL21b]. miscible [Sun22]. Mixed [GL22, PK22, ATW22, BJPV20, CNP22, CLY22, ER20, GQS23, Sun22]. model [BR20, Ber21, BNW20, BHL20, CLY22, CHL21a, Cia22, DES21, FHS20, Huy22]. modelling [BBMRB21]. models [CZ22, OS22]. modes [DBR22]. modified [SS23a]. moment [LL21b]. Monge [Ber20, BySTZ21, GT24]. Monk [CHL21a]. monotone [CMR21]. monotonicity [LZ20]. Monte [LRS20, GV23, GKK^{+}24, LRS19, LSV21, SSW20]. mortar [DS21]. much [Wan21]. multi [Liu21]. multi-dimensional [Liu21]. multicomponent [MNR22]. multidimensional [JNPS23, WS21]. multifidelity [XN23]. Multilevel [CHL^{+}21b, GV23, KW23, LSV21]. multiparameter [RDYL21]. Multipatch [BDK^{+}20]. multiple [DS21, HLS22]. multipoint [CLY22]. multiscale [BJPV20]. multistep [RLK20]. Multivariate [NN21, NP21a].

Navier [AN20, BBMRB21, BD21, FLMS23, Fun22, KN21, LM21, Li21, Mat22, MSSCT21]. NCVEM [CKP22]. near [ABNP21]. Nearest [NP21b]. negative [MMV22]. nematic [BNW20]. Nernst [HH20, SX21, GF22]. nestedness [LL21a]. networks [LLMP24, Uhl24]. Neumann [BRS23, KK20, XY21]. neural [Uhl24]. Newton [Huy22]. node [TNW21]. noise [HJP20, SSW20]. noise-level-robustness [SSW20]. Non [HK22a, BOS22, BK22b, BFM22, CNP22, CMR21, CH20, CH21, Fun22, Huy22, ON22, PR23]. Non-asymptotic [HK22a]. non-autonomous [BOS22]. non-coercive [CMR21]. non-energy [CH20, CH21]. non-linear [BK22b, PR23]. non-monotone [CMR21]. non-selfadjoint [CNP22]. non-stationary [BFM22]. non-stiff [ON22]. non-symmetric [Huy22]. nonequilibrium [MNR22]. nonholonomic [MV20, SMdD22]. Nonlinear [DJC20, CHHM22, HPSV21, Lei22, LW21, TLO22a, TLO22b, XRV21]. nonlocal [AHV24]. Nonnegative [JNPS23]. nonsymmetric [FP22]. norm [DL23, Zen23]. normal [Win20]. norms [CWS22, CWS23, CH20, CH21, MMV22]. note [LO20, LO21]. nuclear [DL23]. number [BCG20, Jun24]. Numerical [ABNP21, BR20, Ber21, CEG^{+}20, CMR21, CZ21, DL23, EGK22, GN21, MT20, NSD20, Pöt23, Zha21, AA21, ALP23, BBFR21, BCD20, BCG20, BD21, CJT21, FCH22a, FCH22b, GSTU20, GSY20, KH20, KN21, Li21, SS22, SNO20, ZJK23]. numerics [TLO22a, TLO22b]. NURBS [DS21].

oblique [CCDS23]. obstacle [Füh20, HKP20]. one [Cia22]. only [ABMS21]. open [AA21]. Operator [BGJR23, AK21, CWHPV23, DST22, DES21, ES24, GSUT21, MCT20, SNO20]. operators [AK20, BAMR22, CEG^{+}20, CEP21, DS22, GSH23, HPSV21, SS23b]. Optimal [CMS20, GSUT21, XL20, AN20, BBFR21, BGIP23, Ber20, BS21, BMS23b, CP24, CJT21, CP23, DP20, DS21, ERSS20, FK23, GSY20, GKK^{+}24, HPSV21, HT22, JL20, JWL22, PK22, VZ22]. Optimization [DLO20, BDP22, ES24, GV23, HJK21, MSA23, NP21b]. optimized [CP22, DNT21]. option [Jun24]. order [ADGS23, AP21, AB21, BBMRB21, BPR21, CCZ20, CEP21, CT22, CKP22, CNP22, CNRS23, CLX21, DRG^{+}20, FPT20, Füh20, HL21a, HL21b, JLZ20, NN21, ZJN21]. oriented [BGIP23]. orthogonal [Van23, XL20]. Orthogonality [BW21]. orthonormal [CIV21]. Oscillation [KV21]. outflow [KN21]. overlapping [GGS22].

P2D [Ber21]. Panov [GTV22]. Panov-type [GTV22]. Parabolic [GKK^{+}24, AM22, Ber20, BPR21, CHY23, GS24, GLW21, GSY20, KNV21, MP23, SS23b]. ParaDiag [KMZ23]. parallel [EF21, GGS22]. parallel-in-time [EF21]. parameter [GSH23, JWL22]. parameter-dependent [GSH23]. parameters [ABMS21]. parametric [BL24, JZZ22]. Part [CJT21, Cia22]. partially [JWL22]. particle [HLW20, HL20a, HLS22, Le 23, LSV21]. partitions [KKK22b]. patch [BKR21]. path [LCO24]. PDE [GKK^{+}24]. PDE-constrained [GKK^{+}24]. PDEs [CNP22, EF21, FMS21, GS24, HSS^{+}20, NN21, TNW21]. penalized [JWL22]. penalty [BySTZ21]. perfectly [DRG^{+}20]. periodic [BAMR22, CEG^{+}20, HHK^{+}24, KKK^{+}22a, LW21, Zha21]. Perthame [GJT20]. perturbed [CS23b, KH20]. Petrov [PK22]. physical [WS21]. physical-constraint-preserving [WS21]. pipe [GL22]. Pitaevskii [AHP21]. pivot [SS23a]. Planck [GF22, DHL20, HH20, JZZ22, NSD20, SX21]. Plane [IGS21, TLO22a, TLO22b]. plasmas [FRZ21]. plate [FHS20]. plus [BDG20]. PML [CHL21a]. point [HI23, KKK^{+}22a]. points [CP22]. Pointwise [KK20]. Poisson [DFTW20, GF22, HH20, Le 23, SX21]. polar [DJC20]. polygon [CW20]. polygon-based [CW20]. polygonal [BJPV20, BySTZ21]. polygons [BMS23a]. polyhedra [OS22]. polyhedral [CWS22, CWS23]. polynomial [AK22, Loi20, Wan21, XL20]. polynomials [MS20, Van23]. porous [BGM22b, Sun22]. port [BGJR23]. port-Hamiltonian [BGJR23]. posed [CDE20, Jin22, PS22]. positive [FLNS21]. positivity [HH20, SX21]. positivity-preserving [HH20]. possibly [Fun22]. posteriori [AM22, BCGT23, CCZ20, CKP22, CFEV21, GLW21, GÖSS20, Kop20, KV21, KPP^{+}21, LV22, MP21, MP23, SSVS21]. potential [MRT20]. power [DJC20]. power-like [DJC20]. Precision [HI23]. Precision-aware [HI23]. preconditioner [AK20]. preconditioners [AA21]. Preconditioning [AP21, BB22, FLNS21, GSUT21, MCT20, MSSCT21]. predictive [Uhl24]. preliminary [BCD20]. PRESB [AK20]. preserving [BL24, BBJP20, BNW20, CHY23, FRZ21, HH20, LSV21, SX21, WS21]. pressure [BHL20, Mat22]. pricing [Jun24]. principle [LZ20]. prior [RICO23]. priori [CKP22, CNRS23, HL20b, SSVS21]. prisms [CW20]. probabilistic [HI23]. probability [BDP22, LRS19, LRS20, Sag22]. problem [AHP21, AN20, AL20, BB22, CS23a, CS23b, CH20, CP23, DNT21, DFTW20, FP22, Füh20, HKP20, Har21, LY22, CH21]. problems [AA24, AJH21, BOS22, BK22a, BBFR21, BNO20, BNO22, CCZ20, CKP22, CNRS23, CMR21, CLX21, CZB22, DES21, EBL22, EOR22, FK23, GN21, GK20, GLW21, GSUT21, GSY20, Hal21, HJP20, HK22a, HL20b, HLO20, HK22b, HW21b, Jin22, JK22, KK20, MP23, ON22, PS22, PR23, RLK20, RDYL21, Sag22, SSVS21, SSW20, SNO20, SS23b, Uhl24, Win20, XN23, Zha21, HW21a]. processes [CHL^{+}21b]. product [CIV21, RDYL21]. programming [Chk24]. Projection [MMV22, KS20, Mat22, Wan21]. projector [BCD20, KNV21]. projector-splitting [KNV21]. projectors [SMS22]. proof [SNO20]. propagation [BL22, ER20, JK20]. properties [AK20, DP20, IGS21, KNV21, KLMNS22, PS21, RICO23, SMS22]. Provably [ON22, WS21]. Proximal [Zen23, CLL^{+}22]. proximal-gradient [CLL^{+}22]. pseudo [EOR22]. pseudo-spectral [EOR22]. pseudodifferential [GSUT21]. Publisher [FCH22b].

QMC [NN21]. quadratic [CNRS23, MS20, VZ22]. quadrature [APR22, dAFR20, AJH21, CEG^{+}20, JK20, MR21, Rie23, TNW21]. quadrilaterals [ATW22]. quantification [HHK^{+}24, KKK^{+}22a, Sag22]. quantum [BDP22]. quasi [GV23, GKK^{+}24, HPSV21, MT20, Zen23]. quasi-homogeneous [MT20]. quasi-Monte [GV23, GKK^{+}24]. quasi-optimal [HPSV21].

Radau [AM22]. radiation [HK23]. Random [OZ17, AB21, CDE20, HHK^{+}24, KNV21, NN21, OZ23]. Randomized [Tro22, ES24, VZ22]. rank [BDG20, BV22, JNPS23, KNV21, KMZ23, ZJN21]. Rate [SS23a]. rates [BD21, CP24, DS21, GS24, Jin22, MH21, PS22, XL20]. Rational [APR22, BRS23, FCH22a, FCH22b, PS21, TNW21]. Raus [Rea24]. Raviart [BOS22, BW21, Guo20, Ji23]. reaction [JK22]. real [FP22]. recovery [ABMS21, BV22, Sun22]. reduced [DS22, Pag21]. reduction [DES21]. reference [HMW23]. reflections [LLS21]. regime [BNO20, BNO22]. region [MSA23]. regions [DLO20]. regression [LPU23]. regularising [HJP20]. regularities [XL20]. Regularity [CK20, AM22, FS20, JLZ20, LW21]. regularization [MMV22, MH21, MMRV24, PR23, Win20]. regularized [BK22a, BS21, HKP20, HL20b]. regularizers [MMRV24]. Reissner [FHS20]. related [MSSCT21]. relations [BW21, CJT21]. relative [EG23, GH21, LSW22]. relative-error [LSW22]. relativistic [WS21]. relaxation [BCG20, MNR22, RLK20]. Relaxed [DFTW20]. ReLU [LLMP24]. reordering [OZ17, OZ23]. representations [EMPS20]. resampled [BC24]. residual [GÖSS20]. Resolution [LM21]. respect [Nat20a, Nat20b]. results [AA21, BCD20]. Rham [BDK^{+}20]. Riemannian [GN21, MSA23, NP21b]. rigid [DBR22]. Rigorous [TLO22b, TLO22a]. risk [GKK^{+}24]. Ritz [SMS22]. Ritz-type [SMS22]. roadmap [IGS21]. Robin [BDG22, Har21]. Robust [CCZ20, CP22, CFEV21]. robustness [Nat20a, Nat20b, SSW20]. rods [BR20]. rough [MMV22]. rule [Rea24]. rules [APR22, dAFR20, AJH21, WM21]. Runge [MR21].

sampling [LRS19, LRS20]. satisfying [BCG20]. scalar [CMS20, CH20, CH21, GTV22]. scales [HLS22]. scattering [AA24, CCWG^{+}24, CWHMB21, EBL22, LY22, NKL22, Zha21]. Schatten [Zen23]. scheme [BGM22a, BBJP20, BKR21, BDG22, CHY23, CCDS23, CGT20, CZ22, CFS20, ES24, FLMM20, GJT20, GTV22, GSTU20, HL21a, HL21b, HH20, KNV21, LM21, Nab21, XRV21]. schemes [AB21, BBMRB21, BPR21, BFM22, CT22, CHHM22, CMS20, CJT21, FPT20, FLMS23, FRZ21, GF22, JK22, Kri20, Kri21, KLMNS22, LSV21, RLK20, SX21]. Schrödinger [CEG^{+}20, LW21]. Schur [DBR22, SNO20]. Schwarz [CP22]. screens [CCWG^{+}24, CWHMB21]. SDE [MP21]. Second [BBMRB21, dAFR20, BPR21, CKP22, CNP22, CNRS23, CWS22, CWS23, HL21a, HL21b, JLZ20, LL21b]. second-kind [CWS22, CWS23]. second-moment [LL21b]. Second-order [BBMRB21, CKP22, CNP22, CNRS23, HL21a, HL21b, JLZ20]. self [BKR21]. self-adjoint [BKR21]. selfadjoint [CNP22]. semi [CCDS23, SSVS21]. semi-Lagrangian [CCDS23]. semi-linear [SSVS21]. semigroups [AB21]. semilinear [CNRS23, CMR21, HLO20, HL21a, HL21b]. Sensitivity [BV22]. separable [Yse20]. sequel [Kri21]. sequence [BDK^{+}20]. serendipity [ATW22]. series [XY21]. set [Ji23]. SGFEM [CZB22]. shallow [FHN22]. Shape [BDP22, DLO20]. sharp [ABNP21, HW21a, HW21b, LSW22]. shells [FHN22]. sign [Cia22]. sign-changing [Cia22]. Signorini [CH21, CH20]. simulations [DHL20]. single [HMW23]. singular [BK22a]. singularities [FS20, TNW21]. singularly [CS23b, KH20]. Sinkhorn [Ber20]. size [Li21]. sliced [BL22]. small [Li21]. smooth [Fun22, KK20, Li21]. smoothness [SMS20]. Sobolev [CK20, Van23]. Solution [dAFR20, BR20, GJT20, JZ20, KS20, Li21, Yse24]. solutions [AA24, AJH21, BK22a, Fun22, KN21, MT20, SNO20, Yse20]. solver [HPSV21]. solvers [Huy22]. solves [Nat20b, Nat20a]. Solving [MSA23, Loi20]. some [OS22]. SOR [OZ17, OZ23]. SOR-type [OZ17, OZ23]. sound [EBL22]. sound-hard [EBL22]. space [DHW23, FMS21, GS24, NSD20, PS21, Yse24]. space-time [DHW23, GS24, NSD20]. spaces [BW21, BHL22, CK20, ERSS20, LLMP24, MH21, OS22, Rea24]. spacetimes [GL20]. Sparse [CIV21, KW23]. spatial [GJT20]. spatio [CHL^{+}21b]. spatio-temporal [CHL^{+}21b]. species [CFS20]. spectra [BAMR22]. Spectral [EBL22, Sag22, BMS23a, BB22, EOR22, MSSCT21, XRV21]. spectrally [MSSCT21]. spectrum [CWHPV23, DBR22]. speed [BCG20]. sphere [BBC23, HT22]. spherical [BBC23]. spline [SMS20, SMS22]. splines [LL21a]. splits [CGGH24]. splitting [BGJR23, EOR22, ES24, KNV21, VZ22]. square [AK20]. squares [Füh20, FK23, LM21]. Stability [BK22b, BPR21, CNP22, EG23, GT24, KNV21, LCO24, BDG22, DL23, FCH22a, FCH22b, Har21, SS23a]. Stabilization [BC24, BCGT23]. Stabilization-free [BCGT23]. stabilized [BNO20, BNO22, JK22]. Stable [CZB22, KL17, DNT21, DRG^{+}20, NP21b, NKL21]. staggered [MSSCT21]. state [GSY20]. states [LLW24, NSD20]. stationary [BFM22]. Statistical [RICO23]. Steklov [AL20]. stepping [BFM22, JLZ20, MP21]. stepsize [HLS22]. Stieltjes [APR22]. stiff [HLO20, ON22]. Stochastic [GSH23, ABNP21, BD21, DHW23, EMPS20, ES24]. Stokes [AN20, AK21, BBMRB21, BB22, BD21, BFM22, CLY22, FLMS23, Fun22, GK20, KN21, LM21, Li21, Mat22, MSSCT21]. Stokes-Brinkman [BBMRB21]. Strang [EOR22]. stress [CLY22, HHMX21]. stress-flux [CLY22]. strong [HLW20, HL20a]. strongly [FRZ21]. structure [BL24, BBJP20, BNW20, BDG22, HMW23, SS22]. structure-preserving [BL24, BBJP20, BNW20]. structured [BGW23]. Subdiffusion [JLZ20, JZ20]. subject [BNO20, BNO22]. sublinear [CIV21]. sublinear-time [CIV21]. submanifolds [LRS19, LRS20]. subproblems [MSA23]. sum [WZL22]. summation [HI23]. super [FHKP24]. super-localized [FHKP24]. superconvergence [MR21]. superconvergent [HSS^{+}20]. Superior [AK20]. superstable [ALP23]. supporting [DLO20]. Surface [Guo20, BL24, CDE20, CLL^{+}22, FMS21]. surfaces [BK22b, EGK22, KLL21]. symmetric [FLNS21, HHMX21, Huy22]. system [BCG20, CFS20, DHW23, FLMM20, FLMS23, GF22, NSD20]. systems [BGJR23, BGW23, FLNS21, FMS21, MP21, Pag21, XRV21].

T [LL21a]. T-splines [LL21a]. tangential [BGW23, HL22]. Taylor [DST22]. technique [SS23a, Sun22]. techniques [BB22]. telegraph [KH20]. temporal [CHL^{+}21b]. tensor [BNW20, DL23, DJC20, EMPS20, HHMX21, JNPS23, ZJN21]. term [CIV21, HL20a, JZZ22]. terms [Jun24]. tetrahedral [KKK22b]. thawed [BL22]. their [IGS21, PS21]. theoretic [MSSCT21]. theoretical [BCD20]. theory [LL21b]. thermal [MNR22]. thin [NKL22]. third [ZJN21]. third-order [ZJN21]. Thomas [BW21]. Three [ZZ22, AN20, CGGH24, CZB22, WM21]. three-dimensional [AN20, CGGH24, CZB22]. Time [NKL22, ADGS23, BL22, BFM22, CS23a, CHHM22, CIV21, DHW23, DHL20, EF21, GS24, GÖSS20, HJK21, HLS22, HL21a, HL21b, Huy22, JLZ20, Loi20, MS20, MP21, NSD20, TLO22b, Uhl24, VZ22, TLO22a]. Time-dependent [NKL22, BFM22, CS23a, JLZ20]. time-discrete [HJK21]. time-domain [GÖSS20]. time-sliced [BL22]. time-splitting [VZ22]. time-stepping [BFM22, MP21]. time-varying [Uhl24]. timestepping [GLW21]. tomography [GH21]. topology [BDP22]. torsion [BR20]. total [BK22a, GSTU20, Mat22]. total-variation [BK22a]. trace [BRS23]. traces [BDK^{+}20, WZL22, Yse20]. trajectories [ALP23]. transforms [CIV21]. transmission [Har21]. Transparent [JK20, KH20, EOR22]. transport [BBMRB21, Ber20, BS21, BMS23b, EG23, ERSS20, HT22, JL20]. trapezoidal [WM21]. treatment [CP22]. trees [JK20]. triangles [AP21]. tridiagonal [FP22]. triple [FP22]. trust [MSA23]. Two [CS23a, LLMP24, MCT20, AK20, BCG20, CH20, CH21, FLMM20, FMS21, GH21, GF22, Nat20a, Nat20b, NSD20]. two-by-two [AK20]. two-dimensional [CH20, CH21, GH21]. Two-layer [LLMP24]. Two-level [MCT20, Nat20a, Nat20b]. two-space [FMS21]. two-speed [BCG20]. type [AK21, BK22b, EF21, FMP22, GTV22, HK22b, Lei22, OZ17, OZ23, SMS22, ZJK23].

ultraweak [FHS20]. unbounded [KH20]. Uncertainty [HHK^{+}24, GV23, GKK^{+}24, KKK^{+}22a, Sag22]. Unconditionally [SX21]. unfitted [CLX21, SSVS21]. uniaxially [BNW20]. Unified [CNRS23, Lei22]. uniform [NN21]. unifying [BGW23]. Uniqueness [Har21]. unitary [BDG20]. Universal [BAMR22]. updates [KMZ23]. using [AB21, AK22, GKK^{+}24, HHK^{+}24, HW21a, HW21b, MSA23, NN21, SNO20]. Uzawa [AK21].

validation [MT20]. value [GN21, KK20, MP23, RLK20]. valued [DES21]. values [BB22]. variable [CMS20]. variables [HHK^{+}24, KPW20]. variation [BK22a, GSTU20]. Variational [BCD20, AJH21, BOS22, CGT20, HL20a, Win20]. variations [BC24]. varying [Uhl24]. velocities [FLMM20]. velocity [HL22]. version [MCT20]. versions [ADGS23]. very [Yse24]. veto [HJP20]. via [AM22, BRS23, BKR21, Chk24, EG23, KMZ23, KPW20, KN21, NP21b]. virtual [FMS21]. viscoelasticity [ALP23]. Vlasov [JZZ22, Le 23]. Voigt [AN20]. Volterra [CHY23]. volume [AHV24, CGT20, CZ22, CFS20, CHHM22, FLMM20, FLMS23, GF22, GL20, Nab21].

Wasserstein [CGT20]. water [BBMRB21]. Wave [AA24, BCD20, CJT21, ER20, EF21, HL21a, HL21b, HK22b, JK20, Lei22, LY22, MR21, MRT20, PK22, Sei22]. wave-type [EF21, HK22b, Lei22]. waveguides [Zha21]. wavelet [LPU23]. Waves [IGS21]. weak [Hal21, KS20, KN21, MP21, ON22]. weakly [KLMNS22]. weighted [KPW20, Le 23, OS22]. Whittle [CK20]. Willmore [HL22, KLL21]. windowed [LLW24]. without [HJP20, SNO20]. work [MV20].

Zhang [Uhl24]. zone [CEG^{+}20].

References

Alouges:2021:NPL

[AA21] François Alouges and Martin Averseng. New preconditioners for the Laplace and Helmholtz integral equations on open curves: analytical framework and numerical results. Numerische Mathematik, 148(2):255–292, June 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01189-5;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01189-5.pdf.

Alves:2024:WSP

[AA24] Carlos J. S. Alves and Pedro R. S. Antunes. Wave scattering problems in exterior domains with the method of fundamental solutions. Numerische Mathematik, 156(2):375–394, April 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-024-01395-x.

Alfonsi:2021:GCH

[AB21] Aurélien Alfonsi and Vlad Bally. A generic construction for high order approximation schemes of semigroups using random grids. Numerische Mathematik, 148(4):743–793, August 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01219-2.

Ammari:2021:DLI

[ABMS21] Habib Ammari, Elie Bretin, Pierre Millien, and Laurent Seppecher. A direct linear inversion for discontinuous elastic parameters recovery from internal displacement information only. Numerische Mathematik, 147(1):189–226, January 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-020-01164-6.

Antonopoulou:2021:NAS

[ABNP21] Dimitra Antonopoulou, Lubomír Banas, Robert Nürnberg, and Andreas Prohl. Numerical approximation of the stochastic Cahn–Hilliard equation near the sharp interface limit. Numerische Mathematik, 147(3):505–551, March 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01179-7;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01179-7.pdf.

Aimi:2023:HOT

[ADGS23] Alessandra Aimi, Giulia Di Credico, Heiko Gimperlein, and Ernst P. Stephan. Higher-order time domain boundary elements for elastodynamics: graded meshes and hp versions. Numerische Mathematik, 154(1–2):35–101, June 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-023-01355-x.

Altmann:2021:MGP

[AHP21] Robert Altmann, Patrick Henning, and Daniel Peterseim. The J-method for the Gross–Pitaevskii eigenvalue problem. Numerische Mathematik, 148(3):575–610, July 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01216-5;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01216-5.pdf.

Aggarwal:2024:AFV

[AHV24] Aekta Aggarwal, Helge Holden, and Ganesh Vaidya. On the accuracy of the finite volume approximations to nonlocal conservation laws. Numerische Mathematik, 156(1):237–271, ???? 2024. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01388-2.

Aylwin:2021:EQR

[AJH21] Rubén Aylwin and Carlos Jerez-Hanckes. The effect of quadrature rules on finite element solutions of Maxwell variational problems. Numerische Mathematik, 147(4):903–936, April 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01186-8.

Axelsson:2020:SPP

[AK20] Owe Axelsson and János Karátson. Superior properties of the PRESB preconditioner for operators on two-by-two block form with square blocks. Numerische Mathematik, 146(2):335–368, October 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01143-x;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01143-x.pdf.

Axelsson:2021:KIU

[AK21] Owe Axelsson and János Karátson. Krylov improvements of the Uzawa method for Stokes type operator matrices. Numerische Mathematik, 148(3):611–631, July 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01208-5;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01208-5.pdf.

Allen:2022:BCP

[AK22] Larry Allen and Robert C. Kirby. Bounds-constrained polynomial approximation using the Bernstein basis. Numerische Mathematik, 152(1):101–126, September 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-022-01311-1.

Armentano:2020:SEP

[AL20] María G. Armentano and Ariel L. Lombardi. The Steklov eigenvalue problem in a cuspidal domain. Numerische Mathematik, 144(2):237–270, February 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-019-01092-0.

Antonietti:2023:LST

[ALP23] Paola F. Antonietti, Lorenzo Liverani, and Vittorino Pata. Lack of superstable trajectories in linear viscoelasticity: a numerical approach. Numerische Mathematik, 153(4):611–633, April 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-023-01351-1.

Akrivis:2022:PEE

[AM22] Georgios Akrivis and Charalambos G. Makridakis. A posteriori error estimates for Radau IIA methods via maximal parabolic regularity. Numerische Mathematik, 150(3):691–717, March 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-022-01271-6.

Anh:2020:DGA

[AN20] Cung The Anh and Tran Minh Nguyet. Discontinuous Galerkin approximations for an optimal control problem of threedimensional Navier–Stokes–Voigt equations. Numerische Mathematik, 145(4):727–769, August 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-020-01132-0.

Ainsworth:2021:PHO

[AP21] Mark Ainsworth and Charles Parker. Preconditioning high order H^{2} conforming finite elements on triangles. Numerische Mathematik, 148(2):223–254, June 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-021-01206-7.

Alahmadi:2022:RGQ

[APR22] J. Alahmadi, M. Pranić, and L. Reichel. Rational Gauss quadrature rules for the approximation of matrix functionals involving Stieltjes functions. Numerische Mathematik, 151(2):443–473, June 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-022-01293-0.

Arbogast:2022:DSM

[ATW22] Todd Arbogast, Zhen Tao, and Chuning Wang. Direct serendipity and mixed finite elements on convex quadrilaterals. Numerische Mathematik, 150(4):929–974, April 2022. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-022-01274-3.

Ben-Artzi:2022:UAC

[BAMR22] Jonathan Ben-Artzi, Marco Marletta, and Frank Rösler. Universal algorithms for computing spectra of periodic operators. Numerische Mathematik, 150(3):719–767, March 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
021-01265-w.

Beik:2022:PTC

[BB22] Fatemeh Panjeh Ali Beik and Michele Benzi. Preconditioning techniques for the coupled Stokes–Darcy problem: spectral and field-of-values analysis. Numerische Mathematik, 150(2):257–298, February 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01267-8;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01267-8.pdf.

Bellet:2023:ICS

[BBC23] Jean-Baptiste Bellet, Matthieu Brachet, and Jean-Pierre Croisille. Interpolation on the cubed sphere with spherical harmonics. Numerische Mathematik, 153(2–3):249–278, March 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01340-w.

Bauer:2022:KIA

[BBF22] M. Bauer, M. Bebendorf, and B. Feist. Kernel-independent adaptive construction of H^{2}-matrix approximations. Numerische Mathematik, 150(1):1–32, January 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01255-y;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01255-y.pdf.

Bartsch:2021:NIB

[BBFR21] Jan Bartsch, Alfio Borzì, Francesco Fanelli, and Souvik Roy. A numerical investigation of Brockett’s ensemble optimal control problems. Numerische Mathematik, 149(1):1–42, September 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01223-6;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01223-6.pdf.

Bonizzoni:2020:SPD

[BBJP20] Francesca Bonizzoni, Marcel Braukhoff, Ansgar Jüngel, and Ilaria Perugia. A structure-preserving discontinuous Galerkin scheme for the Fisher–KPP equation. Numerische Mathematik, 146(1):119–157, September 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01136-w;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01136-w.pdf.

Baird:2021:SOS

[BBMRB21] Graham Baird, Raimund Bürger, Paul E. Méndez, and Ricardo Ruiz-Baier. Second-order schemes for axisymmetric Navier– Stokes-Brinkman and transport equations modelling water filters. Numerische Mathematik, 147(2):431–479, February 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-020-01169-1.

Barbarino:2024:SVA

[BC24] Giovanni Barbarino and Antonio Cicone. Stabilization and variations to the adaptive local iterative filtering algorithm: the fast resampled iterative filtering method. Numerische Mathematik, 156 (2):395–433, April 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-024-01394-y.

Blanc:2020:VPA

[BCD20] X. Blanc, E. Cancès, and M.-S. Dupuy. Variational projector augmented-wave method: theoretical analysis and preliminary numerical results. Numerische Mathematik, 144(2):271– 321, February 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01082-2.

Bouchut:2020:EST

[BCG20] François Bouchut, Christophe Chalons, and Sébastien Guisset. An entropy satisfying two-speed relaxation system for the barotropic Euler equations: application to the numerical approximation of low Mach number flows. Numerische Mathematik, 145(1):35–76, May 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01111-5.

Bertrand:2023:SFH

[BCGT23] Fleurianne Bertrand, Carsten Carstensen, Benedikt Gräßle, and Ngoc Tien Tran. Stabilization-free HHO a posteriori error control. Numerische Mathematik, 154(3–4):369–408, August 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-023-01366-8.

Breit:2021:CRN

[BD21] Dominic Breit and Alan Dodgson. Convergence rates for the numerical approximation of the 2D stochastic Navier–Stokes equations. Numerische Mathematik, 147(3):553–578, March 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01181-z;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01181-z.pdf.

Bevilacqua:2020:FIU

[BDG20] Roberto Bevilacqua, Gianna M. Del Corso, and Luca Gemignani. Fast QR iterations for unitary plus low rank matrices. Numerische Mathematik, 144(1):23–53, January 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01080-4.

Burman:2022:FDL

[BDG22] Erik Burman, Rebecca Durst, and Johnny Guzmán. Fully discrete loosely coupled Robin–Robin scheme for incompressible fluidstructure interaction: stability and error analysis. Numerische Mathematik, 151(4):807–840, August 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-022-01295-y.

Buffa:2020:MAR

[BDK^{+}20] Annalisa Buffa, Jürgen Dölz, Stefan Kurz, Sebastian Schöps, Rafael Vázquez, and Felix Wolf. Multipatch approximation of the de Rham sequence and its traces in isogeometric analysis.

Numerische Mathematik, 144(1):201–236, January 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01079-x.

Braida:2022:STO

[BDP22] B. Braida, J. Dalphin, and Y. Privat. Shape and topology optimization for maximum probability domains in quantum chemistry. Numerische Mathematik, 151(4):1017–1064, August 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-022-01305-z.

Berman:2020:SAP

[Ber20] Robert J. Berman. The Sinkhorn algorithm, parabolic optimal transport and geometric Monge–Ampère equations. Numerische Mathematik, 145(4):771–836, August 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01127-x;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01127-x.pdf.

Bermejo:2021:NAF

[Ber21] R. Bermejo. Numerical analysis of a finite element formulation of the P2D model for lithium-ion cells. Numerische Mathematik, 149(3):463–505, November 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01235-2;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01235-2.pdf.

Burman:2022:ETS

[BFM22] Erik Burman, Stefan Frei, and Andre Massing. Eulerian time-stepping schemes for the non-stationary Stokes equations on time-dependent domains. Numerische Mathematik, 150(2):423–478, February 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01264-x;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01264-x.pdf.

Becker:2023:GOA

[BGIP23] Roland Becker, Gregor Gantner, Michael Innerberger, and Dirk Praetorius. Goal-oriented adaptive finite element methods with optimal computational complexity. Numerische Mathematik, 153 (1):111–140, January 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01334-8.

Bartel:2023:OSB

[BGJR23] Andreas Bartel, Michael Günther, Birgit Jacob, and Timo Reis. Operator splitting based dynamic iteration for linear differentialalgebraic port-Hamiltonian systems. Numerische Mathematik, 155(1–2):1–34, October 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-023-01369-5.

Bellotti:2022:FDF

[BGM22a] Thomas Bellotti, Benjamin Graille, and Marc Massot. Finite difference formulation of any lattice Boltzmann scheme. Numerische Mathematik, 152(1):1–40, September 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-022-01302-2.

Bessaih:2022:CDA

[BGM22b] H. Bessaih, V. Ginting, and B. McCaskill. Continuous data assimilation for displacement in a porous medium. Numerische Mathematik, 151(4):927–962, August 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-022-01306-y.

Benner:2023:UFT

[BGW23] Peter Benner, Serkan Gugercin, and Steffen W. R. Werner. A unifying framework for tangential interpolation of structured bilinear control systems. Numerische Mathematik, 155(3–4):445–483, December 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-023-01380-w.

Burman:2020:CFE

[BHL20] Erik Burman, Peter Hansbo, and Mats G. Larson. A cut finite element method for a model of pressure in fractured media. Numerische Mathematik, 146(4):783–818, December 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01157-5;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01157-5.pdf.

Burman:2022:CBE

[BHL22] Erik Burman, Peter Hansbo, and Mats G. Larson. CutFEM based on extended finite element spaces. Numerische Mathematik, 152 (2):331–369, October 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01313-z.

Barrenechea:2020:MHM

[BJPV20] Gabriel R. Barrenechea, Fabrice Jaillet, Diego Paredes, and Frédéric Valentin. The multiscale hybrid mixed method in general polygonal meshes. Numerische Mathematik, 145(1):197–237, May 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-020-01103-5.

Bartels:2022:EET

[BK22a] Sören Bartels and Alex Kaltenbach. Error estimates for totalvariation regularized minimization problems with singular dual solutions. Numerische Mathematik, 152(4):881–906, December 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-022-01324-w.

Beschle:2022:SEE

[BK22b] Cedric Aaron Beschle and Balázs Kovács. Stability and error estimates for non-linear Cahn–Hilliard-type equations on evolving surfaces. Numerische Mathematik, 151(1):1–48, May 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01280-5.

Bunder:2021:EFP

[BKR21] J. E. Bunder, I. G. Kevrekidis, and A. J. Roberts. Equation-free patch scheme for efficient computational homogenisation via selfadjoint coupling. Numerische Mathematik, 149(2):229–272, October 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-021-01232-5.

Bergold:2022:EBT

[BL22] Paul Bergold and Caroline Lasser. An error bound for the timesliced thawed Gaussian propagation method. Numerische Mathematik, 152(3):511–551, November 2022. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-022-01319-7.

Bao:2024:SPP

[BL24] Weizhu Bao and Yifei Li. A structure-preserving parametric finite element method for geometric flows with anisotropic surface energy. Numerische Mathematik, 156(2):609–639, April 2024. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
024-01398-8.

Banjai:2023:ECH

[BMS23a] Lehel Banjai, Jens M. Melenk, and Christoph Schwab. Exponential convergence of hp FEM for spectral fractional diffusion in polygons. Numerische Mathematik, 153(1):1–47, January 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-022-01329-5.

Bonafini:2023:AAD

[BMS23b] Mauro Bonafini, Ismael Medina, and Bernhard Schmitzer. Asymptotic analysis of domain decomposition for optimal transport. Numerische Mathematik, 153(2–3):451–492, March 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-023-01347-x.

Burman:2020:SFE

[BNO20] Erik Burman, Mihai Nechita, and Lauri Oksanen. A stabilized finite element method for inverse problems subject to the convection-diffusion equation. I: Diffusion-dominated regime. Numerische Mathematik, 144(3):451–477, March 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01087-x;   https://link.springer.com/content/pdf/
10.1007/s00211-019-01087-x.pdf.

Burman:2022:SFE

[BNO22] Erik Burman, Mihai Nechita, and Lauri Oksanen. A stabilized finite element method for inverse problems subject to the convection–diffusion equation. II: convection-dominated regime. Numerische Mathematik, 150(3):769–801, March 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01268-1.

Borthagaray:2020:SPF

[BNW20] Juan Pablo Borthagaray, Ricardo H. Nochetto, and Shawn W. Walker. A structure-preserving FEM for the uniaxially constrained Q-tensor model of nematic liquid crystals. Numerische Mathematik, 145(4):837–881, August 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01133-z.

Balci:2022:CRF

[BOS22] Anna Kh. Balci, Christoph Ortner, and Johannes Storn. Crouzeix–Raviart finite element method for non-autonomous variational problems with Lavrentiev gap. Numerische Mathematik, 151(4):779–805, August 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-022-01303-1.

Bokanowski:2021:SCS

[BPR21] Olivier Bokanowski, Athena Picarelli, and Christoph Reisinger. Stability and convergence of second order backward differentiation schemes for parabolic Hamilton–Jacobi–Bellman equations. Numerische Mathematik, 148(1):187–222, May 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01202-x;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01202-x.pdf.

Bartels:2020:NSB

[BR20] Sören Bartels and Philipp Reiter. Numerical solution of a bending-torsion model for elastic rods. Numerische Mathematik, 146(4):661–697, December 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01156-6;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01156-6.pdf.

Benzi:2023:CNE

[BRS23] Michele Benzi, Michele Rinelli, and Igor Simunec. Computation of the von Neumann entropy of large matrices via trace estimators and rational Krylov methods. Numerische Mathematik, 155(3–4): 377–414, December 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-023-01368-6.

Bonafini:2021:DDE

[BS21] Mauro Bonafini and Bernhard Schmitzer. Domain decomposition for entropy regularized optimal transport. Numerische Mathematik, 149(4):819–870, December 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01245-0;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01245-0.pdf.

Biswas:2022:EED

[BTU22] A. Biswas, J. Tian, and S. Ulusoy. Error estimates for deep learning methods in fluid dynamics. Numerische Mathematik, 151 (3):753–777, July 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01294-z.

Breiding:2022:SLR

[BV22] Paul Breiding and Nick Vannieuwenhoven. Sensitivity of low-rank matrix recovery. Numerische Mathematik, 152(4):725–759, December 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-022-01327-7.

Bartels:2021:ORC

[BW21] Sören Bartels and Zhangxian Wang. Orthogonality relations of Crouzeix–Raviart and Raviart–Thomas finite element spaces. Numerische Mathematik, 148(1):127–139, May 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01199-3;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01199-3.pdf.

Brenner:2021:CEC

[BySTZ21] Susanne C. Brenner, Li yeng Sung, Zhiyu Tan, and Hongchao Zhang. A convexity enforcing C^{0} interior penalty method for the Monge–Ampère equation on convex polygonal domains. Numerische Mathematik, 148(3):497–524, July 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01210-x;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01210-x.pdf.

Calzola:2023:SLS

[CCDS23] Elisa Calzola, Elisabetta Carlini, Xavier Dupuis, and Francisco J. Silva. A semi-Lagrangian scheme for Hamilton–Jacobi–Bellman equations with oblique derivatives boundary conditions. Numerische Mathematik, 153(1):49–84, January 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01336-6.

Caetano:2024:HMB

[CCWG^{+}24] A. M. Caetano, S. N. Chandler-Wilde, A. Gibbs, D. P. Hewett, and A. Moiola. A Hausdorff-measure boundary element method for acoustic scattering by fractal screens. Numerische Mathematik, 156(2):463–532, April 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-024-01399-7.

Cai:2020:REP

[CCZ20] Difeng Cai, Zhiqiang Cai, and Shun Zhang. Robust equilibrated a posteriori error estimator for higher order finite element approximations to diffusion problems. Numerische Mathematik, 144 (1):1–21, January 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01075-1.

Church:2020:DMA

[CDE20] Lewis Church, Ana Djurdjevac, and Charles M. Elliott. A domain mapping approach for elliptic equations posed on random bulk and surface domains. Numerische Mathematik, 146(1):1–49, September 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01139-7;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01139-7.pdf.

Cances:2020:NQB

[CEG^{+}20] Éric Cancès, Virginie Ehrlacher, David Gontier, Antoine Levitt, and Damiano Lombardi. Numerical quadrature in the Brillouin zone for periodic Schrödinger operators. Numerische Mathematik, 144(3):479–526, March 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-019-01096-w.

Carstensen:2021:GLB

[CEP21] Carsten Carstensen, Alexandre Ern, and Sophie Puttkammer. Guaranteed lower bounds on eigenvalues of elliptic operators with a hybrid high-order method. Numerische Mathematik, 149(2):273–304, October 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01228-1;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01228-1.pdf.

Chaumont-Frelet:2021:DGR

[CFEV21] T. Chaumont-Frelet, A. Ern, and M. Vohralík. On the derivation of guaranteed and p-robust a posteriori error estimates for the Helmholtz equation. Numerische Mathematik, 148(3):525–573, July 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-021-01192-w.

Carrillo:2020:CFV

[CFS20] José A. Carrillo, Francis Filbet, and Markus Schmidtchen. Convergence of a finite volume scheme for a system of interacting species with cross-diffusion. Numerische Mathematik, 145(3):473–511, July 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01121-3;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01121-3.pdf.

Christiansen:2024:DEC

[CGGH24] Snorre H. Christiansen, Jay Gopalakrishnan, Johnny Guzmán, and Kaibo Hu. A discrete elasticity complex on three-dimensional Alfeld splits. Numerische Mathematik, 156(1):159–204, ???? 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-023-01381-9.

Cances:2020:VFV

[CGT20] Clément Cancès, Thomas O. Gallouët, and Gabriele Todeschi. A variational finite volume scheme for Wasserstein gradient flows. Numerische Mathematik, 146(3):437–480, November 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01153-9.

Christof:2020:FEE

[CH20] Constantin Christof and Christof Haubner. Finite element error estimates in non-energy norms for the two-dimensional scalar Signorini problem. Numerische Mathematik, 145(3):513–551, July 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-020-01117-z. See correction [CH21].

Christof:2021:CFE

[CH21] Constantin Christof and Christof Haubner. Correction to: Finite element error estimates in non-energy norms for the two-dimensional scalar Signorini problem. Numerische Mathematik, 148(2):495, June 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01204-9;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01204-9.pdf. See [CH20].

Chainais-Hillairet:2022:LTB

[CHHM22] Claire Chainais-Hillairet, Maxime Herda, and Julien Moatti. Long-time behaviour of hybrid finite volume schemes for advection–diffusion equations: linear and nonlinear approaches. Numerische Mathematik, 151(4):963–1016, August 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01289-w.

Chkifa:2024:LEL

[Chk24] Moulay Abdellah Chkifa. Lattice enumeration via linear programming. Numerische Mathematik, 156(1):71–106, ???? 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-023-01376-6.

Chen:2021:DAN

[CHL21a] Meng Chen, Yunqing Huang, and Jichun Li. Development and analysis of a new finite element method for the Cohen–Monk PML model. Numerische Mathematik, 147(1):127–155, January 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-020-01166-4.

Chernov:2021:MEK

[CHL^{+}21b] Alexey Chernov, Håkon Hoel, Kody J. H. Law, Fabio Nobile, and Raul Tempone. Multilevel ensemble Kalman filtering for spatio-temporal processes. Numerische Mathematik, 147(1):71–125, January 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-020-01159-3;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01159-3.pdf.

Calvez:2023:CLV

[CHY23] Vincent Calvez, Hélène Hivert, and Havva Yoldas. Concentration in Lotka–Volterra parabolic equations: an asymptotic-preserving scheme. Numerische Mathematik, 154(1–2):103–153, June 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-023-01362-y.

Ciarlet:2022:AEF

[Cia22] Patrick Ciarlet, Jr. On the approximation of electromagnetic fields by edge finite elements — Part 4: analysis of the model with one sign-changing coefficient. Numerische Mathematik, 152 (2):223–257, October 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01315-x.

Choi:2021:SHT

[CIV21] Bosu Choi, Mark Iwen, and Toni Volkmer. Sparse harmonic transforms II: best s-term approximation guarantees for bounded orthonormal product bases in sublinear-time. Numerische Mathematik, 148(2):293–362, June 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-021-01200-z.

Chen:2021:CNS

[CJT21] Qingshan Chen, Lili Ju, and Roger Temam. Conservative numerical schemes with optimal dispersive wave relations: Part I. Derivation and analysis. Numerische Mathematik, 149(1): 43–85, September 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01218-3.

Cox:2020:RCA

[CK20] Sonja G. Cox and Kristin Kirchner. Regularity and convergence analysis in Sobolev and Hölder spaces for generalized Whittle– Matérn fields. Numerische Mathematik, 146(4):819–873, December 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-020-01151-x;   https://link.springer.com/
content/pdf/10.1007/s00211-020-01151-x.pdf.

Carstensen:2022:PPE

[CKP22] Carsten Carstensen, Rekha Khot, and Amiya K. Pani. A priori and a posteriori error analysis of the lowest-order NCVEM for second-order linear indefinite elliptic problems. Numerische Mathematik, 151(3):551–600, July 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-022-01296-x.

Craig:2022:PGA

[CLL^{+}22] Katy Craig, Jian-Guo Liu, Jianfeng Lu, Jeremy L. Marzuola, and Li Wang. A proximal-gradient algorithm for crystal surface evolution. Numerische Mathematik, 152(3):631–662, November 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-022-01320-0.

Chen:2021:AHO

[CLX21] Zhiming Chen, Ke Li, and Xueshuang Xiang. An adaptive highorder unfitted finite element method for elliptic interface problems. Numerische Mathematik, 149(3):507–548, November 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01243-2.

Caucao:2022:MSF

[CLY22] Sergio Caucao, Tongtong Li, and Ivan Yotov. A multipoint stressflux mixed finite element method for the Stokes–Biot model. Numerische Mathematik, 152(2):411–473, October 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01310-2.

Casas:2021:NAC

[CMR21] Eduardo Casas, Mariano Mateos, and Arnd Rösch. Numerical approximation of control problems of non-monotone and non-coercive semilinear elliptic equations. Numerische Mathematik, 149(2):305–340, October 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01222-7;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01222-7.pdf.

Chen:2020:OEE

[CMS20] Hongtao Chen, Jingjing Mao, and Jie Shen. Optimal error estimates for the scalar auxiliary variable finite-element schemes for gradient flows. Numerische Mathematik, 145(1):167–196, May 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-020-01112-4.

Carstensen:2022:SMF

[CNP22] C. Carstensen, Neela Nataraj, and Amiya K. Pani. Stability of mixed FEMs for non-selfadjoint indefinite second-order linear elliptic PDEs. Numerische Mathematik, 150(4):975–992, April 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-022-01282-3.

Carstensen:2023:UPA

[CNRS23] Carsten Carstensen, Neela Nataraj, Gopikrishnan C. Remesan, and Devika Shylaja. Unified a priori analysis of four secondorder FEM for fourth-order quadratic semilinear problems. Numerische Mathematik, 154(3–4):323–368, August 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01356-w.

Claeys:2022:RTC

[CP22] Xavier Claeys and Emile Parolin. Robust treatment of crosspoints in optimized Schwarz methods. Numerische Mathematik, 151(2):405–442, June 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01288-x.

Chrysafinos:2023:AAO

[CP23] Konstantinos Chrysafinos and Dimitra Plaka. Analysis and approximations of an optimal control problem for the Allen– Cahn equation. Numerische Mathematik, 155(1–2):35–82, October 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-023-01374-8.

Carstensen:2024:AGL

[CP24] Carsten Carstensen and Sophie Puttkammer. Adaptive guaranteed lower eigenvalue bounds with optimal convergence rates. Numerische Mathematik, 156(1):1–38, ???? 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01382-8.

Curry:2020:CLG

[CS20] Charles Curry and Alexander Schmeding. Convergence of Lie group integrators. Numerische Mathematik, 144(2):357– 373, February 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01083-1.

Carstensen:2023:TDT

[CS23a] C. Carstensen and M. Schedensack. Two discretisations of the time-dependent Bingham problem. Numerische Mathematik, 153 (2–3):411–450, March 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01338-4.

Cheng:2023:LDG

[CS23b] Yao Cheng and Martin Stynes. The local discontinuous Galerkin method for a singularly perturbed convection–diffusion problem with characteristic and exponential layers. Numerische Mathematik, 154(1–2):283–318, June 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-023-01361-z.

Carstensen:2022:CAH

[CT22] Carsten Carstensen and Ngoc Tien Tran. Convergent adaptive hybrid higher-order schemes for convex minimization. Numerische Mathematik, 151(2):329–367, June 2022. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-022-01284-1.

Chen:2020:CEP

[CW20] Wenbin Chen and Yanqiu Wang. H^{1}, H(curl) and H(div) conforming elements on polygon-based prisms and cones. Numerische Mathematik, 145(4):973–1004, August 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01129-9.

Chandler-Wilde:2021:BEM

[CWHMB21] Simon N. Chandler-Wilde, David P. Hewett, Andrea Moiola, and Jeanne Besson. Boundary element methods for acoustic scattering by fractal screens. Numerische Mathematik, 147(4):785–837, April 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01182-y;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01182-y.pdf.

Chandler-Wilde:2023:SDL

[CWHPV23] Simon N. Chandler-Wilde, Raffael Hagger, Karl-Mikael Perfekt, and Jani A. Virtanen. On the spectrum of the double-layer operator on locally-dilation-invariant Lipschitz domains. Numerische Mathematik, 153(4):635–699, April 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-023-01353-z.

Chandler-Wilde:2022:CEN

[CWS22] S. N. Chandler-Wilde and E. A. Spence. Coercivity, essential norms, and the Galerkin method for second-kind integral equations on polyhedral and Lipschitz domains. Numerische Mathematik, 150(2):299–371, February 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01256-x;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01256-x.pdf. See correction [CWS23].

Chandler-Wilde:2023:CCE

[CWS23] S. N. Chandler-Wilde and E. A. Spence. Correction to: Coercivity, essential norms, and the Galerkin method for secondkind integral equations on polyhedral and Lipschitz domains. Numerische Mathematik, 154(1–2):319–321, June 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-

023-01357-9. See [CWS22].

Cicone:2021:NAI

[CZ21] Antonio Cicone and Haomin Zhou. Numerical analysis for iterative filtering with new efficient implementations based on FFT. Numerische Mathematik, 147(1):1–28, January 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-020-01165-5.

Cances:2022:CFV

[CZ22] Clément Cancès and Antoine Zurek. A convergent finite volume scheme for dissipation driven models with volume filling constraint. Numerische Mathematik, 151(1):279–328, May 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01270-7.

Cui:2022:SGF

[CZB22] Cu Cui, Qinghui Zhang, and Ivo Babuska. Stable generalized finite element method (SGFEM) for three-dimensional crack problems. Numerische Mathematik, 152(2):475–509, October 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-022-01312-0.

Alba:2020:SSK

[dAFR20] Patricia Díaz de Alba, Luisa Fermo, and Giuseppe Rodriguez. Solution of second kind Fredholm integral equations by means of Gauss and anti-Gauss quadrature rules. Numerische Mathematik, 146(4):699–728, December 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01163-7;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01163-7.pdf.

Dostal:2022:SSC

[DBR22] Zdenek Dostál, Tomás Brzobohatý, and Lubomír Ríha. On the spectrum of Schur complements of 2D elastic clusters joined by rigid edge modes and hybrid domain decomposition. Numerische Mathematik, 152(1):41–66, September 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-022-01307-x.

Dolz:2021:MRA

[DES21] Jürgen Dölz, Herbert Egger, and Matthias Schlottbom. A model reduction approach for inverse problems with operator valued data. Numerische Mathematik, 148(4):889–917, August 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01224-5;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01224-5.pdf.

Diening:2020:RKI

[DFTW20] L. Diening, M. Fornasier, R. Tomasi, and M. Wank. A relaxed Kacanov iteration for the p-Poisson problem. Numerische Mathematik, 145(1):1–34, May 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01107-1;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01107-1.pdf.

Dujardin:2020:CHE

[DHL20] Guillaume Dujardin, Frédéric Hérau, and Pauline Lafitte. Coercivity, hypocoercivity, exponential time decay and simulations for discrete Fokker–Planck equations. Numerische Mathematik, 144 (3):615–697, March 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01094-y.

Diening:2023:AST

[DHW23] Lars Diening, Martina Hofmanová, and Jörn Wichmann. An averaged space-time discretization of the stochastic p-Laplace system. Numerische Mathematik, 153(2–3):557–609, March 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-022-01343-7.

Dong:2020:NPL

[DJC20] Bo Dong, Nan Jiang, and Moody T. Chu. Nonlinear power-like iteration by polar decomposition and its application to tensor approximation. Numerische Mathematik, 144(4):729–749, April 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-020-01100-8.

Dai:2023:NST

[DL23] Zhen Dai and Lek-Heng Lim. Numerical stability and tensor nuclear norm. Numerische Mathematik, 155(3–4):345–376, December 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-023-01377-5.

Dapogny:2020:OSR

[DLO20] Charles Dapogny, Nicolas Lebbe, and Edouard Oudet. Optimization of the shape of regions supporting boundary conditions. Numerische Mathematik, 146(1):51–104, September 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01140-0.

Despres:2021:CSO

[DNT21] B. Després, A. Nicolopoulos, and B. Thierry. Corners and stable optimized domain decomposition methods for the Helmholtz problem. Numerische Mathematik, 149(4):779–818, December 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-021-01251-2.

Delgado:2020:EOP

[DP20] Jorge Delgado and J. M. Peña. Extremal and optimal properties of B-bases collocation matrices. Numerische Mathematik, 146(1): 105–118, September 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01135-x.

Duru:2020:SDG

[DRG^{+}20] Kenneth Duru, Leonhard Rannabauer, Alice-Agnes Gabriel, Gunilla Kreiss, and Michael Bader. A stable discontinuous Galerkin method for the perfectly matched layer for elastodynamics in first order form. Numerische Mathematik, 146(4):729–782, December 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-020-01160-w.

Dornisch:2021:IMM

[DS21] W. Dornisch and J. Stöckler. An isogeometric mortar method for the coupling of multiple NURBS domains with optimal convergence rates. Numerische Mathematik, 149(4):871–931, December 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01246-z;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01246-z.pdf.

Danczul:2022:RBM

[DS22] Tobias Danczul and Joachim Schöberl. A reduced basis method for fractional diffusion operators I. Numerische Mathematik, 151 (2):369–404, June 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01287-y.

Diening:2022:FOT

[DST22] L. Diening, J. Storn, and T. Tscherpel. Fortin operator for the Taylor–Hood element. Numerische Mathematik, 150 (2):671–689, February 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01260-1;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01260-1.pdf.

Ecevit:2022:SGB

[EBL22] Fatih Ecevit, Yassine Boubendir, and Souaad Lazergui. Spectral Galerkin boundary element methods for high-frequency soundhard scattering problems. Numerische Mathematik, 150(3): 803–847, March 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01269-0.

Ellison:2021:PTA

[EF21] Abe C. Ellison and Bengt Fornberg. A parallel-in-time approach for wave-type PDEs. Numerische Mathematik, 148(1):79–98, May 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-021-01197-5.

Egger:2023:SAA

[EG23] H. Egger and J. Giesselmann. Stability and asymptotic analysis for instationary gas transport via relative energy estimates. Numerische Mathematik, 153(4):701–728, April 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01349-9.

Elliott:2022:NAI

[EGK22] Charles M. Elliott, Harald Garcke, and Balázs Kovács. Numerical analysis for the interaction of mean curvature flow and diffusion on closed surfaces. Numerische Mathematik, 151(4):873–925, August 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-022-01301-3.

Eigel:2020:ASG

[EMPS20] Martin Eigel, Manuel Marschall, Max Pfeffer, and Reinhold Schneider. Adaptive stochastic Galerkin FEM for lognormal coefficients in hierarchical tensor representations. Numerische Mathematik, 145(3):655–692, July 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01123-1;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01123-1.pdf.

Einkemmer:2022:PSS

[EOR22] L. Einkemmer, A. Ostermann, and M. Residori. A pseudospectral Strang splitting method for linear dispersive problems with transparent boundary conditions. Numerische Mathematik, 150(1):105–135, January 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01252-1;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01252-1.pdf.

Egger:2020:MLM

[ER20] H. Egger and B. Radu. A mass-lumped mixed finite element method for acoustic wave propagation. Numerische Mathematik, 145(2):239–269, June 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01118-y.

Erbar:2020:COT

[ERSS20] Matthias Erbar, Martin Rumpf, Bernhard Schmitzer, and Stefan Simon. Computation of optimal transport on discrete metric measure spaces. Numerische Mathematik, 144(1):157–200, January 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-019-01077-z.

Eisenmann:2024:ROS

[ES24] Monika Eisenmann and Tony Stillfjord. A randomized operator splitting scheme inspired by stochastic optimization methods. Numerische Mathematik, 156(2):435–461, April 2024. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
024-01396-w.

Fuda:2022:NSL

[FCH22a] Chiara Fuda, Rosanna Campagna, and Kai Hormann. On the numerical stability of linear barycentric rational interpolation. Numerische Mathematik, 152(4):761–786, December 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-

022-01316-w. See publisher correction [FCH22b].

Fuda:2022:PCN

[FCH22b] Chiara Fuda, Rosanna Campagna, and Kai Hormann. Publisher correction to: On the numerical stability of linear barycentric rational interpolation. Numerische Mathematik, 152(4):787– 788, December 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01330-y. See [FCH22a].

Freese:2024:SLG

[FHKP24] Philip Freese, Moritz Hauck, Tim Keil, and Daniel Peterseim. A super-localized generalized finite element method. Numerische Mathematik, 156(1):205–235, ???? 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-023-01386-4.

Fuhrer:2022:DMS

[FHN22] Thomas Führer, Norbert Heuer, and Antti H. Niemi. A DPG method for shallow shells. Numerische Mathematik, 152(1):67– 99, September 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01308-w.

Fuhrer:2020:UFR

[FHS20] Thomas Führer, Norbert Heuer, and Francisco-Javier Sayas. An ultraweak formulation of the Reissner–Mindlin plate bending model and DPG approximation. Numerische Mathematik, 145 (2):313–344, June 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01116-0.

Fuhrer:2023:LSF

[FK23] Thomas Führer and Michael Karkulik. Least-squares finite elements for distributed optimal control problems. Numerische Mathematik, 154(3–4):409–442, August 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01367-7.

Feireisl:2020:FVS

[FLMM20] Eduard Feireisl, Mária Lukácová-Medvid’ová, and Hana Mizerová. A finite volume scheme for the Euler system inspired by the two velocities approach. Numerische Mathematik, 144(1): 89–132, January 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01078-y.

Feireisl:2023:IEE

[FLMS23] Eduard Feireisl, Mária Lukácová-Medvid’ová, and Bangwei She. Improved error estimates for the finite volume and the MAC schemes for the compressible Navier–Stokes system. Numerische Mathematik, 153(2–3):493–529, March 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-023-01346-y.

Fang:2021:PSP

[FLNS21] Zhi-Wei Fang, Xue-Lei Lin, Michael K. Ng, and Hai-Wei Sun. Preconditioning for symmetric positive definite systems in balanced fractional diffusion equations. Numerische Mathematik, 147 (3):651–677, March 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01175-x.

Faustmann:2022:CTE

[FMP22] Markus Faustmann, Jens Markus Melenk, and Maryam Parvizi. Caccioppoli-type estimates and H-matrix approximations to inverses for FEM-BEM couplings. Numerische Mathematik, 150 (3):849–892, March 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-021-01261-0.

Frittelli:2021:BSV

[FMS21] Massimo Frittelli, Anotida Madzvamuse, and Ivonne Sgura. Bulksurface virtual element method for systems of PDEs in two-space dimensions. Numerische Mathematik, 147(2):305–348, February 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-020-01167-3;   https://link.springer.com/
content/pdf/10.1007/s00211-020-01167-3.pdf.

Ferreira:2022:RTD

[FP22] Carla Ferreira and Beresford Parlett. A real triple dqds algorithm for the nonsymmetric tridiagonal eigenvalue problem. Numerische Mathematik, 150(2):373–422, February 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01254-z.

Falcone:2020:CAF

[FPT20] Maurizio Falcone, Giulio Paolucci, and Silvia Tozza. Convergence of adaptive filtered schemes for first order evolutionary Hamilton– Jacobi equations. Numerische Mathematik, 145(2):271–311, June 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-020-01119-x.

Filbet:2021:CAA

[FRZ21] Francis Filbet, L. Miguel Rodrigues, and Hamed Zakerzadeh. Convergence analysis of asymptotic preserving schemes for strongly magnetized plasmas. Numerische Mathematik, 149(3): 549–593, November 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01248-x.

Feischl:2020:ECH

[FS20] M. Feischl and Ch. Schwab. Exponential convergence in H^{1} of hp-FEM for Gevrey regularity with isotropic singularities. Numerische Mathematik, 144(2):323–346, February 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01085-z;   https://link.springer.com/content/pdf/
10.1007/s00211-019-01085-z.pdf.

Fuhrer:2020:FOL

[Füh20] Thomas Führer. First-order least-squares method for the obstacle problem. Numerische Mathematik, 144(1):55–88, January 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-019-01084-0.

Funaro:2022:HWN

[Fun22] Daniele Funaro. How and why non smooth solutions of the 3D Navier–Stokes equations could possibly develop. Numerische Mathematik, 152(4):789–817, December 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01333-9.

Gaudeul:2022:ECA

[GF22] Benoît Gaudeul and Jürgen Fuhrmann. Entropy and convergence analysis for two finite volume schemes for a Nernst–Planck– Poisson system with ion volume constraints. Numerische Mathematik, 151(1):99–149, May 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-022-01279-y.

Gong:2022:CPO

[GGS22] Shihua Gong, Martin J. Gander, and Euan A. Spence. Convergence of parallel overlapping domain decomposition methods for the Helmholtz equation. Numerische Mathematik, 152(2): 259–306, October 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01318-8.

Garde:2021:MRC

[GH21] Henrik Garde and Nuutti Hyvönen. Mimicking relative continuum measurements by electrode data in two-dimensional electrical impedance tomography. Numerische Mathematik, 147(3):579–609, March 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-020-01170-8;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01170-8.pdf.

Ghoshal:2020:CGS

[GJT20] Shyam Sundar Ghoshal, Animesh Jana, and John D. Towers. Convergence of a Godunov scheme to an Audusse–Perthame adapted entropy solution for conservation laws with BV spatial flux. Numerische Mathematik, 146(3):629–659, November 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-020-01150-y.

Gedicke:2020:DCD

[GK20] Joscha Gedicke and Arbaz Khan. Divergence-conforming discontinuous Galerkin finite elements for Stokes eigenvalue problems. Numerische Mathematik, 144(3):585–614, March 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01095-x;   https://link.springer.com/content/pdf/
10.1007/s00211-019-01095-x.pdf.

Guth:2024:PPC

[GKK^{+}24] Philipp A. Guth, Vesa Kaarnioja, Frances Y. Kuo, Claudia Schillings, and Ian H. Sloan. Parabolic PDE-constrained optimal control under uncertainty with entropic risk measure using quasi-Monte Carlo integration. Numerische Mathematik, 156 (2):565–608, April 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-024-01397-9.

Giesselmann:2020:FCF

[GL20] Jan Giesselmann and Philippe G. LeFloch. Formulation and convergence of the finite volume method for conservation laws on spacetimes with boundary. Numerische Mathematik, 144(4): 751–785, April 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01101-7.

Gustafsson:2022:MFE

[GL22] Tom Gustafsson and Philip L. Lederer. Mixed finite elements for Bingham flow in a pipe. Numerische Mathematik, 152(4): 819–840, December 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01332-w.

Georgoulis:2021:PEB

[GLW21] Emmanuil H. Georgoulis, Omar Lakkis, and Thomas P. Wihler. A posteriori error bounds for fully-discrete hp-discontinuous Galerkin timestepping methods for parabolic problems. Numerische Mathematik, 148(2):363–386, June 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01187-7;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01187-7.pdf.

Garcke:2021:NAB

[GN21] Harald Garcke and Robert Nürnberg. Numerical approximation of boundary value problems for curvature flow and elastic flow in Riemannian manifolds. Numerische Mathematik, 149(2):375–415, October 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01231-6;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01231-6.pdf.

Gimperlein:2020:RPE

[GÖSS20] Heiko Gimperlein, Ceyhun Özdemir, David Stark, and Ernst P. Stephan. A residual a posteriori error estimate for the timedomain boundary element method. Numerische Mathematik, 146 (2):239–280, October 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01142-y.

Gao:2023:NAM

[GQS23] Huadong Gao, Weifeng Qiu, and Weiwei Sun. New analysis of mixed FEMs for dynamical incompressible magnetohydrodynamics. Numerische Mathematik, 153(2–3):327–358, March 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01341-9.

Gantner:2024:IRS

[GS24] Gregor Gantner and Rob Stevenson. Improved rates for a spacetime FOSLS of parabolic PDEs. Numerische Mathematik, 156 (1):133–157, ???? 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-023-01387-3.

Grubisic:2023:SCM

[GSH23] Luka Grubisić, Mikael Saarikangas, and Harri Hakula. Stochastic collocation method for computing eigenspaces of parameterdependent operators. Numerische Mathematik, 153(1):85–110, January 2023. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL https://link.springer.com/
article/10.1007/s00211-022-01339-3.

Giga:2020:NNS

[GSTU20] Yoshikazu Giga, Koya Sakakibara, Kazutoshi Taguchi, and Masaaki Uesaka. A new numerical scheme for discrete constrained total variation flows and its convergence. Numerische Mathematik, 146(1):181–217, September 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01134-y;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01134-y.pdf.

Gimperlein:2021:OOP

[GSUT21] Heiko Gimperlein, Jakub Stocek, and Carolina Urzúa-Torres. Optimal operator preconditioning for pseudodifferential boundary problems. Numerische Mathematik, 148(1):1–41, May 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01193-9.

Glowinski:2020:ANA

[GSY20] Roland Glowinski, Yongcun Song, and Xiaoming Yuan. An ADMM numerical approach to linear parabolic state constrained optimal control problems. Numerische Mathematik, 144(4): 931–966, April 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01104-4.

Gallistl:2024:SGE

[GT24] Dietmar Gallistl and Ngoc Tien Tran. Stability and guaranteed error control of approximations to the Monge–Ampère equation. Numerische Mathematik, 156(1):107–131, ???? 2024. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01385-5.

Ghoshal:2022:GTS

[GTV22] Shyam Sundar Ghoshal, John D. Towers, and Ganesh Vaidya. A Godunov type scheme and error estimates for scalar conservation laws with Panov-type discontinuous flux. Numerische Mathematik, 151(3):601–625, July 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-022-01297-w.

Guo:2020:SCR

[Guo20] Hailong Guo. Surface Crouzeix–Raviart element for the Laplace– Beltrami equation. Numerische Mathematik, 144(3):527–551, March 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-019-01099-7.

Guth:2023:MQM

[GV23] Philipp A. Guth and Andreas Van Barel. Multilevel quasi-Monte Carlo for optimization under uncertainty. Numerische Mathematik, 154(3–4):443–484, August 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-023-01364-w.

Halla:2021:GAH

[Hal21] Martin Halla. Galerkin approximation of holomorphic eigenvalue problems: weak T -coercivity and T -compatibility. Numerische Mathematik, 148(2):387–407, June 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01205-8;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01205-8.pdf.

Harrach:2021:USG

[Har21] Bastian Harrach. Uniqueness, stability and global convergence for a discrete inverse elliptic Robin transmission problem. Numerische Mathematik, 147(1):29–70, January 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
020-01162-8;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01162-8.pdf.

Hu:2020:FDP

[HH20] Jingwei Hu and Xiaodong Huang. A fully discrete positivitypreserving and energy-dissipative finite difference scheme for Poisson–Nernst–Planck equations. Numerische Mathematik, 145 (1):77–115, May 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01109-z.

Hakula:2024:UQR

[HHK^{+}24] Harri Hakula, Helmut Harbrecht, Vesa Kaarnioja, Frances Y. Kuo, and Ian H. Sloan. Uncertainty quantification for random domains using periodic random variables. Numerische Mathematik, 156(1):273–317, ???? 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-023-01392-6.

Hong:2021:NDG

[HHMX21] Qingguo Hong, Jun Hu, Limin Ma, and Jinchao Xu. New discontinuous Galerkin algorithms and analysis for linear elasticity with symmetric stress tensor. Numerische Mathematik, 149(3): 645–678, November 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01234-3.

Hallman:2023:PAD

[HI23] Eric Hallman and Ilse C. F. Ipsen. Precision-aware deterministic and probabilistic error bounds for floating point summation. Numerische Mathematik, 155(1–2):83–119, October 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01370-y.

Ha:2021:CEE

[HJK21] Seung-Yeal Ha, Shi Jin, and Doheon Kim. Convergence and error estimates for time-discrete consensus-based optimization algorithms. Numerische Mathematik, 147(2):255–282, February 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01174-y.

Harrach:2020:BBV

[HJP20] Bastian Harrach, Tim Jahn, and Roland Potthast. Beyond the Bakushinkii veto: regularising linear inverse problems without knowing the noise distribution. Numerische Mathematik, 145(3):581–603, July 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01122-2;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01122-2.pdf.

Helin:2022:NAE

[HK22a] Tapio Helin and Remo Kretschmann. Non-asymptotic error estimates for the Laplace approximation in Bayesian inverse problems. Numerische Mathematik, 150(2):521–549, February 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01266-9;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01266-9.pdf.

Hochbruck:2022:EAF

[HK22b] Marlis Hochbruck and Jonas Köhler. Error analysis of a fully discrete discontinuous Galerkin alternating direction implicit discretization of a class of linear wave-type problems. Numerische Mathematik, 150(3):893–927, March 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-021-01262-z.

Hagstrom:2023:CRB

[HK23] Thomas Hagstrom and Seungil Kim. Complete radiation boundary conditions for the Helmholtz equation II: domains with corners. Numerische Mathematik, 153(4):775–825, April 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01352-0.

Hafemeyer:2020:FEE

[HKP20] Dominik Hafemeyer, Christian Kahle, and Johannes Pfefferer. Finite element error estimates in L^{2} for regularized discrete approximations to the obstacle problem. Numerische Mathematik, 144 (1):133–156, January 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01081-3.

Hairer:2020:LTA

[HL20a] Ernst Hairer and Christian Lubich. Long-term analysis of a variational integrator for charged-particle dynamics in a strong magnetic field. Numerische Mathematik, 144(3):699–728, March 2020.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-019-01093-z.

Heltai:2020:PEE

[HL20b] Luca Heltai and Wenyu Lei. A priori error estimates of regularized elliptic problems. Numerische Mathematik, 146(3):571– 596, November 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01152-w.

Hochbruck:2021:CIE

[HL21a] Marlis Hochbruck and Jan Leibold. Correction to: An implicitexplicit time discretization scheme for second-order semilinear wave equations with application to dynamic boundary conditions. Numerische Mathematik, 147(4):901, April 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01190-y;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01190-y.pdf. See [HL21b].

Hochbruck:2021:IET

[HL21b] Marlis Hochbruck and Jan Leibold. An implicit-explicit time discretization scheme for second-order semilinear wave equations with application to dynamic boundary conditions. Numerische Mathematik, 147(4):869–899, April 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01184-w;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01184-w.pdf. See correction [HL21a].

Hu:2022:EFE

[HL22] Jiashun Hu and Buyang Li. Evolving finite element methods with an artificial tangential velocity for mean curvature flow and Willmore flow. Numerische Mathematik, 152(1):127–181, September 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-022-01309-9.

Hochbruck:2020:CLM

[HLO20] Marlis Hochbruck, Jan Leibold, and Alexander Ostermann. On the convergence of Lawson methods for semilinear stiff problems. Numerische Mathematik, 145(3):553–580, July 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01120-4;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01120-4.pdf.

Hairer:2022:LSI

[HLS22] Ernst Hairer, Christian Lubich, and Yanyan Shi. Large-stepsize integrators for charged-particle dynamics over multiple time scales. Numerische Mathematik, 151(3):659–691, July 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01298-9.

Hairer:2020:FBA

[HLW20] Ernst Hairer, Christian Lubich, and Bin Wang. A filtered Boris algorithm for charged-particle dynamics in a strong magnetic field. Numerische Mathematik, 144(4):787–809, April 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01105-3;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01105-3.pdf.

Hassan:2023:ASR

[HMW23] Muhammad Hassan, Yvon Maday, and Yipeng Wang. Analysis of the single reference coupled cluster method for electronic structure calculations: the full-coupled cluster equations. Numerische Mathematik, 155(1–2):121–173, October 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01371-x.

Haberl:2021:CQO

[HPSV21] Alexander Haberl, Dirk Praetorius, Stefan Schimanko, and Martin Vohralík. Convergence and quasi-optimal cost of adaptive algorithms for nonlinear operators including iterative linearization and algebraic solver. Numerische Mathematik, 147(3):679–725, March 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01176-w;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01176-w.pdf.

Hu:2020:SHD

[HSS^{+}20] Weiwei Hu, Jiguang Shen, John R. Singler, Yangwen Zhang, and Xiaobo Zheng. A superconvergent hybridizable discontinuous Galerkin method for Dirichlet boundary control of elliptic PDEs. Numerische Mathematik, 144(2):375–411, February 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-019-01090-2.

Hamfeldt:2022:CFO

[HT22] Brittany Froese Hamfeldt and Axel G. R. Turnquist. A convergence framework for optimal transport on the sphere. Numerische Mathematik, 151(3):627–657, July 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-022-01292-1.

Huynh:2022:NKB

[Huy22] Ngoc Mai Monica Huynh. Newton–Krylov-BDDC deluxe solvers for non-symmetric fully implicit time discretizations of the bidomain model. Numerische Mathematik, 152(4):841–879, December 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-022-01331-x.

Hollbacher:2021:CSI

[HW21a] Susanne Höllbacher and Gabriel Wittum. Correction to: A sharp interface method using enriched finite elements for elliptic interface problems. Numerische Mathematik, 147(4):783, April 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01191-x;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01191-x.pdf. See [HW21b].

Hollbacher:2021:SIM

[HW21b] Susanne Höllbacher and Gabriel Wittum. A sharp interface method using enriched finite elements for elliptic interface problems. Numerische Mathematik, 147(4):759–781, April 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01180-0;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01180-0.pdf. See correction [HW21a].

Imbert-Gerard:2021:RGP

[IGS21] Lise-Marie Imbert-Gérard and Guillaume Sylvand. A roadmap for Generalized Plane Waves and their interpolation properties. Numerische Mathematik, 149(1):87–137, September 2021.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01220-9.

Ji:2023:ICR

[Ji23] Haifeng Ji. An immersed Crouzeix–Raviart finite element method in 2D and 3D based on discrete level set functions. Numerische Mathematik, 153(2–3):279–325, March 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-023-01345-z.

Jin:2022:CRD

[Jin22] Qinian Jin. Convergence rates of a dual gradient method for constrained linear ill-posed problems. Numerische Mathematik, 151 (4):841–871, August 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01300-4.

Joly:2020:TBC

[JK20] Patrick Joly and Maryna Kachanovska. Transparent boundary conditions for wave propagation in fractal trees: convolution quadrature approach. Numerische Mathematik, 146(2): 281–334, October 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01145-9.

John:2022:ASS

[JK22] Volker John and Petr Knobloch. On algebraically stabilized schemes for convection-diffusion-reaction problems. Numerische Mathematik, 152(3):553–585, November 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01325-9.

Jacobs:2020:FAO

[JL20] Matt Jacobs and Flavien Léger. A fast approach to optimal transport: the back-and-forth method. Numerische Mathematik, 146(3):513–544, November 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-020-01154-8.

Jin:2020:STD

[JLZ20] Bangti Jin, Buyang Li, and Zhi Zhou. Subdiffusion with timedependent coefficients: improved regularity and second-order time stepping. Numerische Mathematik, 145(4):883–913, August 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-020-01130-2.

Jiang:2023:NLR

[JNPS23] Tai-Xiang Jiang, Michael K. Ng, Junjun Pan, and Guang-Jing Song. Nonnegative low rank tensor approximations with multidimensional image applications. Numerische Mathematik, 153(1): 141–170, January 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01328-6.

Junike:2024:NTC

[Jun24] Gero Junike. On the number of terms in the COS method for European option pricing. Numerische Mathematik, 156(2):533–564, April 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-024-01402-1.

Ji:2022:NPF

[JWL22] Haifeng Ji, Feng Wang, and Zhilin Li. A new parameter free partially penalized immersed finite element and the optimal convergence analysis. Numerische Mathematik, 150(4):1035–1086, April 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-022-01276-1.

Jin:2020:IIS

[JZ20] Bangti Jin and Zhi Zhou. Incomplete iterative solution of subdiffusion. Numerische Mathematik, 145(3):693–725, July 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-020-01128-w.

Jin:2022:VFP

[JZZ22] Shi Jin, Yuhua Zhu, and Enrique Zuazua. The Vlasov–Fokker– Planck equation with high dimensional parametric forcing term. Numerische Mathematik, 150(2):479–519, February 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01257-w.

Kong:2020:TBC

[KH20] Wang Kong and Zhongyi Huang. Transparent boundary conditions and numerical computation for singularly perturbed telegraph equation on unbounded domain. Numerische Mathematik, 145(2):345–382, June 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01115-1.

Kashiwabara:2020:PEE

[KK20] Takahito Kashiwabara and Tomoya Kemmochi. Pointwise error estimates of linear finite element method for Neumann boundary value problems in a smooth domain. Numerische Mathematik, 144 (3):553–584, March 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01098-8.

Kaarnioja:2022:FAP

[KKK^{+}22a] Vesa Kaarnioja, Yoshihito Kazashi, Frances Y. Kuo, Fabio Nobile, and Ian H. Sloan. Fast approximation by periodic kernelbased lattice-point interpolation with application in uncertainty quantification. Numerische Mathematik, 150(1):33–77, January 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01242-3;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01242-3.pdf.

Korotov:2022:DFE

[KKK22b] Sergey Korotov, Michal Krízek, and Václav Kucera. On degenerating finite element tetrahedral partitions. Numerische Mathematik, 152(2):307–329, October 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-022-01317-9.

Kovacs:2017:SCF

[KL17] Balázs Kovács and Christian Lubich. Stable and convergent fully discrete interior–exterior coupling of Maxwell’s equations. Numerische Mathematik, 137(1):91–117, September 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See correction [NKL21].

Kovacs:2021:CEF

[KLL21] Balázs Kovács, Buyang Li, and Christian Lubich. A convergent evolving finite element algorithm for Willmore flow of closed surfaces. Numerische Mathematik, 149(3):595–643, November 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01238-z.

Kucera:2022:APC

[KLMNS22] Václav Kucera, Mária Lukácová-Medvid’ová, Sebastian Noelle, and Jochen Schütz. Asymptotic properties of a class of linearly implicit schemes for weakly compressible Euler equations. Numerische Mathematik, 150(1):79–103, January 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01240-5;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01240-5.pdf.

Kressner:2023:IPL

[KMZ23] Daniel Kressner, Stefano Massei, and Junli Zhu. Improved Para-Diag via low-rank updates and interpolation. Numerische Mathematik, 155(1–2):175–209, October 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-023-01372-w.

Kwon:2021:CWS

[KN21] Young-Sam Kwon and Antonin Novotný. Construction of weak solutions to compressible Navier–Stokes equations with general inflow/outflow boundary conditions via a numerical approximation. Numerische Mathematik, 149(4):717–778, December 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01237-0.

Kazashi:2021:SPP

[KNV21] Yoshihito Kazashi, Fabio Nobile, and Eva Vidlicková. Stability properties of a projector-splitting scheme for dynamical low rank approximation of random parabolic equations. Numerische Mathematik, 149(4):973–1024, December 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01241-4;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01241-4.pdf.

Kopteva:2020:LPE

[Kop20] Natalia Kopteva. Lower a posteriori error estimates on anisotropic meshes. Numerische Mathematik, 146(1):159–179, September 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-020-01137-9.

Kurz:2021:FPE

[KPP^{+}21] Stefan Kurz, Dirk Pauly, Dirk Praetorius, Sergey Repin, and Daniel Sebastian. Functional a posteriori error estimates for boundary element methods. Numerische Mathematik, 147(4):937–966, April 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01188-6;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01188-6.pdf.

Kritzer:2020:EWI

[KPW20] P. Kritzer, F. Pillichshammer, and G. W. Wasilkowski. On efficient weighted integration via a change of variables. Numerische Mathematik, 146(3):545–570, November 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01147-7.

Kriel:2020:EIF

[Kri20] A. J. Kriel. Entropy inequalities for fully-discrete E-schemes. Numerische Mathematik, 144(2):347–356, February 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01089-9.

Kriel:2021:EIF

[Kri21] A. J. Kriel. Entropy inequalities for fully-discrete E-schemes: a sequel. Numerische Mathematik, 149(1):139–149, September 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01227-2.

Kuroki:2020:CCP

[KS20] Hidesato Kuroki and Kohei Soga. On convergence of Chorin’s projection method to a Leray–Hopf weak solution. Numerische Mathematik, 146(2):401–433, October 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01144-w.

Kreuzer:2021:OPE

[KV21] Christian Kreuzer and Andreas Veeser. Oscillation in a posteriori error estimation. Numerische Mathematik, 148(1):43–78, May 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01194-8;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01194-8.pdf.

Kempf:2023:HDA

[KW23] Rüdiger Kempf and Holger Wendland. High-dimensional approximation with kernel-based multilevel methods on sparse grids. Numerische Mathematik, 154(3–4):485–519, August 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01363-x.

Liu:2024:SME

[LCO24] Xuanyu Liu, Huajie Chen, and Christoph Ortner. Stability of the minimum energy path. Numerische Mathematik, 156(1):39–70, ???? 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-023-01391-7.

LeHenaff:2023:GFW

[Le 23] Yoann Le Henaff. Grid-free weighted particle method applied to the Vlasov–Poisson equation. Numerische Mathematik, 155(3–4): 289–344, December 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-023-01378-4.

Leibold:2022:UEA

[Lei22] Jan Leibold. A unified error analysis for nonlinear wave-type equations with application to acoustic boundary conditions. Numerische Mathematik, 152(4):907–936, December 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01326-8.

Li:2021:BNS

[Li21] Buyang Li. A bounded numerical solution with a small mesh size implies existence of a smooth solution to the Navier–Stokes equations. Numerische Mathematik, 147(2):283–304, February 2021.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01172-0.

Liu:2021:ADD

[Liu21] Hailiang Liu. Analysis of direct discontinuous Galerkin methods for multi-dimensional convection–diffusion equations. Numerische Mathematik, 147(4):839–867, April 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01183-x.

Li:2021:SAD

[LL21a] Xiliang Li and Xin Li. AS++ T-splines: arbitrary degree, nestedness and approximation. Numerische Mathematik, 148(4): 795–816, August 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01214-7.

Lin:2021:MCS

[LL21b] Jing Lin and Pierre F. J. Lermusiaux. Minimum-correction second-moment matching: theory, algorithms and applications. Numerische Mathematik, 147(3):611–650, March 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01178-8.

Li:2024:TLN

[LLMP24] Yuanyuan Li, Shuai Lu, Peter Mathé, and Sergei V. Pereverzev. Two-layer networks with the ReLU^{k} activation function: Barron spaces and derivative approximation. Numerische Mathematik, 156(1):319–344, ???? 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-023-01384-6.

Laurent:2021:MR

[LLS21] Philippe Laurent, Guillaume Legendre, and Julien Salomon. On the method of reflections. Numerische Mathematik, 148(2): 449–493, June 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01207-6.

Loring:2024:LWL

[LLW24] Terry A. Loring, Jianfeng Lu, and Alexander B. Watson. Locality of the windowed local density of states. Numerische Mathematik, 156(2):741–775, April 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-024-01400-3.

Lemoine:2021:RIE

[LM21] Jérôme Lemoine and Arnaud Münch. Resolution of the implicit Euler scheme for the Navier–Stokes equation through a leastsquares method. Numerische Mathematik, 147(2):349–391, February 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-021-01171-1.

Lange:2020:NDF

[LO20] Marko Lange and Shin’ichi Oishi. A note on Dekker’s FastTwoSum algorithm. Numerische Mathematik, 145(2):383– 403, June 2020. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL https://link.springer.com/
article/10.1007/s00211-020-01114-2. See correction [LO21].

Lange:2021:CND

[LO21] Marko Lange and Shin’ichi Oishi. Correction to: A note on Dekker’s FastTwoSum algorithm. Numerische Mathematik, 149(1):227–228, September 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01213-8;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01213-8.pdf. See [LO20].

Loisel:2020:EAS

[Loi20] Sébastien Loisel. Efficient algorithms for solving the p-Laplacian in polynomial time. Numerische Mathematik, 146 (2):369–400, October 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01141-z;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01141-z.pdf.

Lippert:2023:FHW

[LPU23] Laura Lippert, Daniel Potts, and Tino Ullrich. Fast hyperbolic wavelet regression meets ANOVA. Numerische Mathematik, 154 (1–2):155–207, June 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-023-01358-8.

Lelievre:2019:HMC

[LRS19] Tony Lelièvre, Mathias Rousset, and Gabriel Stoltz. Hybrid Monte Carlo methods for sampling probability measures on submanifolds. Numerische Mathematik, 143(2):379–421, October 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-019-01056-4. See correction [LRS20].

Lelievre:2020:CHM

[LRS20] Tony Lelièvre, Mathias Rousset, and Gabriel Stoltz. Correction to: Hybrid Monte Carlo methods for sampling probability measures on submanifolds. Numerische Mathematik, 144(2):447–449, February 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01086-y;
 https://link.springer.com/content/pdf/10.1007/s00211-
019-01086-y.pdf. See [LRS19].

Lovbak:2021:MMC

[LSV21] Emil Løvbak, Giovanni Samaey, and Stefan Vandewalle. A multilevel Monte Carlo method for asymptotic-preserving particle schemes in the diffusive limit. Numerische Mathematik, 148 (1):141–186, May 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01201-y.

Lafontaine:2022:SRE

[LSW22] D. Lafontaine, E. A. Spence, and J. Wunsch. A sharp relativeerror bound for the Helmholtz h-FEM at high frequency. Numerische Mathematik, 150(1):137–178, January 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01253-0;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01253-0.pdf.

Liu:2022:FCP

[LV22] Xuefeng Liu and Tomás Vejchodský. Fully computable a posteriori error bounds for eigenfunctions. Numerische Mathematik, 152(1):183–221, September 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-022-01304-0.

Li:2021:FDL

[LW21] Buyang Li and Yifei Wu. A fully discrete low-regularity integrator for the 1D periodic cubic nonlinear Schrödinger equation. Numerische Mathematik, 149(1):151–183, September 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01226-3.

Li:2022:AFE

[LY22] Peijun Li and Xiaokai Yuan. An adaptive finite element DtN method for the elastic wave scattering problem. Numerische Mathematik, 150(4):993–1033, April 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-022-01273-4.

Li:2020:MDM

[LZ20] Hao Li and Xiangxiong Zhang. On the monotonicity and discrete maximum principle of the finite difference implementation of C^{0}–Q^{2} finite element method. Numerische Mathematik, 145 (2):437–472, June 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-020-01110-6.

Matsui:2022:PMN

[Mat22] Kazunori Matsui. A projection method for Navier–Stokes equations with a boundary condition including the total pressure. Numerische Mathematik, 152(3):663–699, November 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01323-x.

Marchand:2020:TLP

[MCT20] Pierre Marchand, Xavier Claeys, and Pierre-Henri Tournier. Twolevel preconditioning for h-version boundary element approximation of hypersingular operator with GenEO. Numerische Mathematik, 146(3):597–628, November 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-020-01149-5.

Miller:2021:MSA

[MH21] Philip Miller and Thorsten Hohage. Maximal spaces for approximation rates in ℓ^{1}-regularization. Numerische Mathematik, 149(2):341–374, October 2021. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01225-4;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01225-4.pdf.

Molinari:2024:IRL

[MMRV24] Cesare Molinari, Mathurin Massias, Lorenzo Rosasco, and Silvia Villa. Iterative regularization for low complexity regularizers. Numerische Mathematik, 156(2):641–689, April 2024. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01390-8.

Millar:2022:PNN

[MMV22] F. Millar, I. Muga, and K. G. Van der Zee. Projection in negative norms and the regularization of rough linear functionals. Numerische Mathematik, 150(4):1087–1121, April 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01278-z.

Marmignon:2022:ERA

[MNR22] Claude Marmignon, Fabio Naddei, and Florent Renac. Energy relaxation approximation for compressible multicomponent flows in thermal nonequilibrium. Numerische Mathematik, 151 (1):151–184, May 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01285-0.

Merle:2021:ATS

[MP21] Fabian Merle and Andreas Prohl. An adaptive time-stepping method based on a posteriori weak error analysis for large SDE systems. Numerische Mathematik, 149(2):417–462, October 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01233-4;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01233-4.pdf.

Merle:2023:PEA

[MP23] Fabian Merle and Andreas Prohl. A posteriori error analysis and adaptivity for high-dimensional elliptic and parabolic boundary value problems. Numerische Mathematik, 153(4):827–884, April 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-023-01350-2.

Melenk:2021:SRK

[MR21] Jens Markus Melenk and Alexander Rieder. On superconvergence of Runge–Kutta convolution quadrature for the wave equation. Numerische Mathematik, 147(1):157–188, January 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
020-01161-9;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01161-9.pdf.

Micu:2020:ACW

[MRT20] Sorin Micu, Ionel Roventa, and Laurentiu Emanuel Temereanca. Approximation of the controls for the wave equation with a potential. Numerische Mathematik, 144(4):835–887, April 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01106-2.

Malyshev:2020:CDC

[MS20] Alexander Malyshev and Miloud Sadkane. Computing the distance to continuous-time instability of quadratic matrix polynomials. Numerische Mathematik, 145(1):149–165, May 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01108-0.

Mor:2023:STR

[MSA23] Uria Mor, Boris Shustin, and Haim Avron. Solving trust region subproblems using Riemannian optimization. Numerische Mathematik, 154(1–2):1–33, June 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-023-01360-0.

Mazza:2021:MTS

[MSSCT21] M. Mazza, M. Semplice, S. Serra-Capizzano, and E. Travaglia. A matrix-theoretic spectral analysis of incompressible Navier– Stokes staggered DG approximations and a related spectrally based preconditioning approach. Numerische Mathematik, 149(4): 933–971, December 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01247-y.

Matsue:2020:NVB

[MT20] Kaname Matsue and Akitoshi Takayasu. Numerical validation of blow-up solutions with quasi-homogeneous compactifications. Numerische Mathematik, 145(3):605–654, July 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01125-z.

Modin:2020:WMN

[MV20] Klas Modin and Olivier Verdier. What makes nonholonomic integrators work? Numerische Mathematik, 145(2):405–435, June 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-020-01126-y;   https://link.springer.com/
content/pdf/10.1007/s00211-020-01126-y.pdf.

Nabet:2021:EEF

[Nab21] Flore Nabet. An error estimate for a finite-volume scheme for the Cahn–Hilliard equation with dynamic boundary conditions. Numerische Mathematik, 149(1):185–226, September 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01230-7.

Nataf:2020:CMA

[Nat20a] Frédéric Nataf. Correction to: Mathematical analysis of robustness of two-level domain decomposition methods with respect to inexact coarse solves. Numerische Mathematik, 146(2):435–436, October 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01146-8;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01146-8.pdf. See [Nat20b].

Nataf:2020:MAR

[Nat20b] Frédéric Nataf. Mathematical analysis of robustness of two-level domain decomposition methods with respect to inexact coarse solves. Numerische Mathematik, 144(4):811–833, April 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-

020-01102-6. See correction [Nat20a].

Nick:2021:CSC

[NKL21] Jörg Nick, Balázs Kovács, and Christian Lubich. Correction to: Stable and convergent fully discrete interior–exterior coupling of Maxwell’s equations. Numerische Mathematik, 147(4):997–1000, April 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01196-6;
 https://www.math.utah.edu/pub/tex/bib/nummath2010.bib;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01196-6.pdf. See [KL17].

Nick:2022:TDE

[NKL22] Jörg Nick, Balázs Kovács, and Christian Lubich. Time-dependent electromagnetic scattering from thin layers. Numerische Mathematik, 150(4):1123–1164, April 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-022-01277-0.

Nguyen:2021:MMD

[NN21] Dong T. P. Nguyen and Dirk Nuyens. MDFEM: Multivariate decomposition finite element method for elliptic PDEs with uniform random diffusion coefficients using higher-order QMC and FEM. Numerische Mathematik, 148(3):633–669, July 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01212-9.

Nasdala:2021:EMA

[NP21a] Robert Nasdala and Daniel Potts. Efficient multivariate approximation on the cube. Numerische Mathematik, 147 (2):393–429, February 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01177-9;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01177-9.pdf.

Noferini:2021:NVS

[NP21b] Vanni Noferini and Federico Poloni. Nearest Ω-stable matrix via Riemannian optimization. Numerische Mathematik, 148(4):817–851, August 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01217-4;

 https://link.springer.com/content/pdf/10.1007/s00211-
021-01217-4.pdf.

Nie:2020:NAS

[NSD20] Daxin Nie, Jing Sun, and Weihua Deng. Numerical algorithm for the space-time fractional Fokker–Planck system with two internal states. Numerische Mathematik, 146(3):481–511, November 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-020-01148-6.

OReilly:2022:PNS

[ON22] Ossian O’Reilly and Jan Nordström. Provably non-stiff implementation of weak coupling conditions for hyperbolic problems. Numerische Mathematik, 150(2):551–589, February 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01263-y;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01263-y.pdf.

Otarola:2022:AAS

[OS22] Enrique Otárola and Abner J. Salgado. On the analysis and approximation of some models of fluids over weighted spaces on convex polyhedra. Numerische Mathematik, 151(1):185–218, May 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-022-01272-5.

Oswald:2017:RRS

[OZ17] Peter Oswald and Weiqi Zhou. Random reordering in SORtype methods. Numerische Mathematik, 135(4):1207–1220, April 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0829-7;   http:
//link.springer.com/article/10.1007/s00211-016-0829-7.

See correction [OZ23].

Oswald:2023:CRR

[OZ23] Peter Oswald and Weiqi Zhou. Correction: Random reordering in SOR-type methods. Numerische Mathematik, 154(3–4):521–525, August 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-023-01365-9.

Pagliantini:2021:DRB

[Pag21] Cecilia Pagliantini. Dynamical reduced basis methods for Hamiltonian systems. Numerische Mathematik, 148(2):409–448, June 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.com/article/10.
1007/s00211-021-01211-w;   https://link.springer.com/
content/pdf/10.1007/s00211-021-01211-w.pdf.

Peralta:2022:MHP

[PK22] Gilbert Peralta and Karl Kunisch. Mixed and hybrid Petrov– Galerkin finite element discretization for optimal control of the wave equation. Numerische Mathematik, 150(2):591–627, February 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-021-01258-9.

Potzsche:2023:NDI

[Pöt23] Christian Pötzsche. Numerical dynamics of integrodifference equations. Numerische Mathematik, 154(1–2):249–281, June 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-023-01354-y.

Pieronek:2023:IRN

[PR23] Lukas Pieronek and Andreas Rieder. On the iterative regularization of non-linear illposed problems in L^{∞}. Numerische Mathematik, 154(1–2):209–247, June 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-023-01359-7.

Pozza:2021:FRK

[PS21] Stefano Pozza and Valeria Simoncini. Functions of rational Krylov space matrices and their decay properties. Numerische Mathematik, 148(1):99–126, May 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-021-01198-4.

Parzer:2022:CRA

[PS22] Fabian Parzer and Otmar Scherzer. On convergence rates of adaptive ensemble Kalman inversion for linear ill-posed problems. Numerische Mathematik, 152(2):371–409, October 2022.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-022-01314-y.

Rodriguez:2021:FPH

[RDYL21] Jose Israel Rodriguez, Jin-Hong Du, Yiling You, and Lek-Heng Lim. Fiber product homotopy method for multiparameter eigenvalue problems. Numerische Mathematik, 148(4):853–888, August 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-021-01215-6.

Real:2024:HRR

[Rea24] Rommel R. Real. Hanke–Raus rule for Landweber iteration in Banach spaces. Numerische Mathematik, 156(1):345–373, ???? 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-023-01389-1.

Reid:2023:SPB

[RICO23] Tim W. Reid, Ilse C. F. Ipsen, Jon Cockayne, and Chris J. Oates. Statistical properties of BayesCG under the Krylov prior. Numerische Mathematik, 155(3–4):239–288, December 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01375-7.

Rieder:2023:DEQ

[Rie23] Alexander Rieder. Double exponential quadrature for fractional diffusion. Numerische Mathematik, 153(2–3):359–410, March 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-022-01342-8.

Ranocha:2020:GRM

[RLK20] Hendrik Ranocha, Lajos Lóczi, and David I. Ketcheson. General relaxation methods for initial-value problems with application to multistep schemes. Numerische Mathematik, 146(4):875–906, December 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-020-01158-4.

Sagiv:2022:SCP

[Sag22] Amir Sagiv. Spectral convergence of probability densities for forward problems in uncertainty quantification. Numerische Mathematik, 150(4):1165–1186, April 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-022-01281-4.

Seibel:2022:BEM

[Sei22] Daniel Seibel. Boundary element methods for the wave equation based on hierarchical matrices and adaptive cross approximation. Numerische Mathematik, 150(2):629–670, February 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
021-01259-8;   https://link.springer.com/content/pdf/
10.1007/s00211-021-01259-8.pdf.

Simoes:2022:EDL

[SMdD22] Alexandre Anahory Simoes, Juan Carlos Marrero, and David Martín de Diego. Exact discrete Lagrangian mechanics for nonholonomic mechanics. Numerische Mathematik, 151(1):49–98, May 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-022-01283-2.

Sande:2020:EEE

[SMS20] Espen Sande, Carla Manni, and Hendrik Speleers. Explicit error estimates for spline approximation of arbitrary smoothness in isogeometric analysis. Numerische Mathematik, 144(4):889–929, April 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-019-01097-9.

Sande:2022:RTP

[SMS22] Espen Sande, Carla Manni, and Hendrik Speleers. Ritz-type projectors with boundary interpolation properties and explicit spline error estimates. Numerische Mathematik, 151(2):475–494, June 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-022-01286-z.

Sekine:2020:NFU

[SNO20] Kouta Sekine, Mitsuhiro T. Nakao, and Shin’ichi Oishi. A new formulation using the Schur complement for the numerical existence proof of solutions to elliptic problems: without direct estimation for an inverse of the linearized operator. Numerische Mathematik, 146(4):907–926, December 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01155-7;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01155-7.pdf.

Schwarzacher:2022:NAF

[SS22] Sebastian Schwarzacher and Bangwei She. On numerical approximations to fluid-structure interactions involving compressible fluids. Numerische Mathematik, 151(1):219–278, May 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
022-01275-2.

Saha:2023:RCS

[SS23a] Jitraj Saha and Mehakpreet Singh. Rate of convergence and stability analysis of a modified fixed pivot technique for a fragmentation equation. Numerische Mathematik, 153(2–3):531–555, March 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-023-01344-0.

Stevenson:2023:IOP

[SS23b] Rob Stevenson and Johannes Storn. Interpolation operators for parabolic problems. Numerische Mathematik, 155(1–2):211– 238, October 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-023-01373-9.

Sanchez:2021:PPE

[SSVS21] Nestor Sánchez, Tonatiuh Sánchez-Vizuet, and Manuel E. Solano. A priori and a posteriori error analysis of an unfitted HDG method for semi-linear elliptic problems. Numerische Mathematik, 148 (4):919–958, August 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01221-8.

Schillings:2020:CLA

[SSW20] Claudia Schillings, Björn Sprungk, and Philipp Wacker. On the convergence of the Laplace approximation and noiselevel-robustness of Laplace-based Monte Carlo methods for Bayesian inverse problems. Numerische Mathematik, 145 (4):915–971, August 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-020-01131-1;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01131-1.pdf.

Sun:2022:NAR

[Sun22] Weiwei Sun. New analysis and recovery technique of mixed FEMs for compressible miscible displacement in porous media. Numerische Mathematik, 150(1):179–215, January 2022. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01249-w.

Shen:2021:UPP

[SX21] Jie Shen and Jie Xu. Unconditionally positivity preserving and energy dissipative schemes for Poisson–Nernst–Planck equations. Numerische Mathematik, 148(3):671–697, July 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01203-w.

Takayasu:2022:CRN

[TLO22a] Akitoshi Takayasu, Jean-Philippe Lessard, and Hisashi Okamoto. Correction to: Rigorous numerics for nonlinear heat equations in the complex plane of time. Numerische Mathematik, 151 (3):751–752, July 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01299-8. See [TLO22b].

Takayasu:2022:RNN

[TLO22b] Akitoshi Takayasu, Jean-Philippe Lessard, and Hisashi Okamoto. Rigorous numerics for nonlinear heat equations in the complex plane of time. Numerische Mathematik, 151(3):693–750, July 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-022-01291-2. See correction [TLO22a].

Trefethen:2021:ENC

[TNW21] Lloyd N. Trefethen, Yuji Nakatsukasa, and J. A. C. Weideman. Exponential node clustering at singularities for rational approximation, quadrature, and PDEs. Numerische Mathematik, 147(1):227–254, January 2021. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-020-01168-2;
 https://link.springer.com/content/pdf/10.1007/s00211-
020-01168-2.pdf.

Tropp:2022:RBK

[Tro22] Joel A. Tropp. Randomized block Krylov methods for approximating extreme eigenvalues. Numerische Mathematik, 150(1): 217–255, January 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01250-3.

Uhlig:2024:ZNN

[Uhl24] Frank Uhlig. Zhang neural networks: an introduction to predictive computations for discretized time-varying matrix problems. Numerische Mathematik, 156(2):691–739, April 2024. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
023-01393-5.

VanBuggenhout:2023:GSO

[Van23] Niel Van Buggenhout. On generating Sobolev orthogonal polynomials. Numerische Mathematik, 155(3–4):415–443, December 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-023-01379-3.

Veldman:2022:FRT

[VZ22] D. W. M. Veldman and E. Zuazua. A framework for randomized time-splitting in linear-quadratic optimal control. Numerische Mathematik, 151(2):495–549, June 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-022-01290-3.

Wang:2021:HMF

[Wan21] Haiyong Wang. How much faster does the best polynomial approximation converge than Legendre projection? Numerische Mathematik, 147(2):481–503, February 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01173-z.

Winkler:2020:EEV

[Win20] Max Winkler. Error estimates for variational normal derivatives and Dirichlet control problems with energy regularization.

Numerische Mathematik, 144(2):413–445, February 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01091-1.

Wu:2021:CTR

[WM21] Bowei Wu and Per-Gunnar Martinsson. Corrected trapezoidal rules for boundary integral equations in three dimensions. Numerische Mathematik, 149(4):1025–1071, December 2021. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01244-1.

Wu:2021:PPC

[WS21] Kailiang Wu and Chi-Wang Shu. Provably physical-constraint-preserving discontinuous Galerkin methods for multidimensional relativistic MHD equations. Numerische Mathematik, 148(3): 699–741, July 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01209-4.

Wang:2022:MSC

[WZL22] Li Wang, Lei-Hong Zhang, and Ren-Cang Li. Maximizing sum of coupled traces with applications. Numerische Mathematik, 152(3):587–629, November 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-022-01322-y.

Xiang:2020:ODR

[XL20] Shuhuang Xiang and Guidong Liu. Optimal decay rates on the asymptotics of orthogonal polynomial expansions for functions of limited regularities. Numerische Mathematik, 145(1):117–148, May 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-020-01113-3.

Xu:2023:BLD

[XN23] Yiming Xu and Akil Narayan. Budget-limited distribution learning in multifidelity problems. Numerische Mathematik, 153(1): 171–212, January 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01337-5.

Xavier:2021:EEF

[XRV21] Juliana C. Xavier, Mauro A. Rincon, and Daniel G. Alfaro Vigo. Error estimates for a fully discrete spectral scheme for nonlinear Boussinesq systems. Numerische Mathematik, 149(3):679– 716, November 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01239-y.

Xu:2021:AFS

[XY21] Liwei Xu and Tao Yin. Analysis of the Fourier series Dirichlet-to-Neumann boundary condition of the Helmholtz equation and its application to finite element methods. Numerische Mathematik, 147(4):967–996, April 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-021-01195-7.

Yserentant:2020:ESL

[Yse20] Harry Yserentant. On the expansion of solutions of Laplace-like equations into traces of separable higher dimensional functions. Numerische Mathematik, 146(1):219–238, September 2020. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
020-01138-8;   https://link.springer.com/content/pdf/
10.1007/s00211-020-01138-8.pdf.

Yserentant:2024:IMS

[Yse24] Harry Yserentant. An iterative method for the solution of Laplace-like equations in high and very high space dimensions. Numerische Mathematik, 156(2):777–811, April 2024. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
024-01401-2.

Zeng:2023:PLM

[Zen23] Chao Zeng. Proximal linearization methods for Schatten pquasi-norm minimization. Numerische Mathematik, 153(1):213– 248, January 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01335-7.

Zhang:2021:NMS

[Zha21] Ruming Zhang. Numerical methods for scattering problems in periodic waveguides. Numerische Mathematik, 148

(4):959–996, August 2021.  CODEN NUMMA7.  ISSN

0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01229-0;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01229-0.pdf.

Zhou:2023:NMC

[ZJK23] Guanyu Zhou, Feifei Jing, and Takahito Kashiwabara. The numerical methods for the coupled fluid flow under the leak interface condition of the friction-type. Numerische Mathematik, 153 (4):729–773, April 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-023-01348-w.

Zeng:2021:AMC

[ZJN21] Chao Zeng, Tai-Xiang Jiang, and Michael K. Ng. An approximation method of CP rank for third-order tensor completion. Numerische Mathematik, 147(3):727–757, March 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-021-01185-9.

Zhang:2022:TFG

[ZZ22] Qian Zhang and Zhimin Zhang. Three families of grad divconforming finite elements. Numerische Mathematik, 152(3):701– 724, November 2022. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-022-01321-z.